### v1

# AI-driven Wireless Positioning: Fundamentals, Standards, State-of-the-art, and Challenges

Abstract
Wireless positioning technologies hold significant value for applications in autonomous driving, extended reality (XR), unmanned aerial vehicles (UAVs), and more. With the advancement of artificial intelligence (AI), leveraging AI to enhance positioning accuracy and robustness has emerged as a field full of potential. Driven by the requirements and functionalities defined in the 3rd Generation Partnership Project (3GPP) standards, AI/machine learning (ML)-based positioning is becoming a key technology to overcome the limitations of traditional methods. This paper begins with an introduction to the fundamentals of AI and wireless positioning, covering AI models, algorithms, positioning applications, emerging wireless technologies, and the basics of positioning techniques. Subsequently, focusing on standardization progress, we provide a comprehensive review of the evolution of 3GPP positioning standards, with an emphasis on the integration of AI/ML technologies in recent and upcoming releases. Based on the AI/ML-assisted positioning and direct AI/ML positioning schemes outlined in the standards, we conduct an in-depth investigation of related research. we focus on state-of-the-art (SOTA) research in AI-based line-of-sight (LOS)/non-line-of-sight (NLOS) detection, time of arrival (TOA)/time difference of arrival (TDOA) estimation, and angle estimation techniques. For Direct AI/ML Positioning, we explore SOTA advancements in fingerprint-based positioning, knowledge-assisted AI positioning, and channel charting-based positioning. Furthermore, we introduce publicly available datasets for wireless positioning and conclude by summarizing the challenges and opportunities of AI-driven wireless positioning.
Index Terms:
Artificial intelligence, positioning technologies, 3GPP, cellular networks, 5G.
I Introduction
With the widespread deployment of 5G networks, wireless positioning technology has become a critical research area. Accurate positioning is indispensable for enabling the effective operation of systems and enhancing user experiences in applications such as intelligent transportation, emergency response, logistics tracking, and extended reality (XR). By providing precise location information, wireless positioning enables more efficient resource management, accurate service delivery, and enhanced security.
Wireless positioning, beyond its communication capabilities, represents a vital application of wireless networks, making its research and performance improvements of significant importance. In outdoor environments, cellular positioning and global navigation satellite system (GNSS) technologies play pivotal roles, providing essential support for pedestrian navigation, autonomous driving, and unmanned aerial vehicle (UAV) localization [1]. These technologies not only enhance positioning accuracy but also improve user experience and safety. Indoors, advancements in technologies such as cellular networks, WiFi, Bluetooth, and ultra-wideband (UWB) have enabled precise localization within complex environments, which is crucial for applications like shopping mall navigation and emergency evacuation [2]. As part of mobile communication infrastructure, cellular positioning is instrumental in both indoor and outdoor scenarios. However, traditional positioning methods, primarily based on geometric relationships like triangulation and trilateration, still face challenges in terms of accuracy, robustness, scalability, and adaptability to dynamic environments.
In recent years, the evolution of artificial intelligence (AI) has brought transformative changes to wireless positioning [3]. AI technologies, particularly machine learning (ML) and deep learning, have significantly enhanced the accuracy and efficiency of positioning systems through their powerful data processing and pattern recognition capabilities. AI algorithms can analyze complex wireless signal characteristics, identify environmental interference factors, and predict user locations with high precision. Furthermore, AI-driven systems can adapt to changing environments by learning from historical data, offering innovative solutions to the challenges of traditional techniques. As AI continues to advance, AI-driven wireless positioning is expected to play an increasingly critical role in the future.
Nevertheless, AI-driven wireless positioning also faces several challenges. First, AI algorithms require substantial amounts of data for training, and obtaining high-quality training data in the wireless positioning domain is often difficult. Additionally, wireless signals are influenced by various factors, such as buildings and weather conditions, line-of-sight (LOS) and non-line-of-sight (NLOS) scenario, which pose challenges to the generalization capabilities of AI algorithms. Moreover, the computational complexity of AI algorithms limits their application in resource-constrained devices. Therefore, designing lightweight, efficient AI algorithms and leveraging limited data resources for effective training are crucial research areas in wireless positioning.
This survey aims to provide a comprehensive overview of AI-driven wireless positioning, encompassing the fundamentals of AI and wireless technologies, progress in standardization, state-of-the-art developments, and associated challenges and research opportunities. Although this paper mainly focuses on AI applications in cellular positioning, the survey also covers some research on AI technology on WiFi, Bluetooth, and UWB systems to present a holistic view of algorithm innovation.
I-A Related Work
Wireless positioning has been a widely researched topic, with extensive investigations focusing on a variety of techniques, scenarios, and applications. Below, we provide a detailed overview of existing surveys and explain how this paper, with its focus on AI-driven wireless positioning, distinguishes itself.
To address the challenges posed by wireless signal positioning, researchers have conducted numerous studies on the fundamental theories of positioning. Focusing on localization algorithms, the authors in [4] provide a comprehensive survey of time-of-arrival (TOA)-based localization algorithms. In [2], the authors summarize indoor positioning algorithms and discuss systems based on WiFi, radio frequency identification (RFID), UWB, Bluetooth, and other technologies. The survey in [5] explores fusion-based indoor positioning techniques using data from cellular networks, WiFi, global positioning systems (GPSs), inertial navigation, cameras, and more. In [6], the authors investigate device-free positioning algorithms. Regarding fingerprint-based positioning technologies, the authors in[7] and [8] conduct detailed surveys, with [9] specifically addressing methods that bypass offline fingerprint maps, and [10] delving deeper into intelligent algorithms for fingerprinting. Considering the impact of network capabilities on positioning technologies, the authors in [11] survey techniques for jammer localization in multi-hop networks, providing critical insights into handling interference and security issues in positioning systems. In [12], the authors explore advanced network localization and tracking technologies. Building on these works, the authors in [13] review the latest research in positioning technologies within wireless networks.
With advancements in technology, positioning capabilities in cellular networks have become increasingly important. In [1], the authors investigate the standardization efforts for cellular positioning from 1G to 4G and summarize the key technologies for wireless positioning in 5G networks. Following this, the paper in [14] provides an in-depth survey of positioning technologies in 5G networks, covering standardization, key elements, research trends, and performance analyses in real-world environments. Looking ahead to 6G networks, the authors in [15] summarize the latest research in positioning technologies, including novel applications, supporting technologies, system models, and critical performance indicators, while also discussing future research directions. Moreover, with advancements in communication technology, D2D networks [16], millimeter-wave (mmWave) [17], THz [18], reconfigurable intelligent surface (RIS) [19], and Ground-Air-Space Networks [20] are not only enhancing communication capabilities but also improving positioning performance. Detailed surveys on these technologically supported positioning systems have been conducted in [16, 17, 18, 19, 20].
The rapid evolution of AI has added a transformative dimension to wireless positioning. ML algorithms, in particular, have been extensively applied to extract features from wireless signals that correlate with environmental factors and user locations, thereby enabling precise and efficient positioning. Most of the above surveys incorporate discussions on ML-based positioning methodologies. Specifically, the authors in [21] provide an in-depth analysis of AI-based indoor positioning technologies. However, these studies did focus on the connection between cellular positioning and AI. Meanwhile, the swift advancements in AI have outpaced the scope of these existing reviews. This gap highlights the pressing need for an updated, comprehensive survey to support current and future research efforts.
I-B Contributions
Figure 1: The overall structure of this paper.
As mentioned above, the rapid advancements in AI technology and the significant progress in AI-driven wireless positioning have motivated this comprehensive survey. This paper aims to systematically explore the fundamental principles of AI and wireless positioning technologies, ongoing standardization efforts, cutting-edge advancements, and the persistent challenges shaping future research directions. Specifically, the main contributions of this paper are as follows:
  • Foundation of AI and Wireless Positioning: This paper provides a comprehensive review of the foundational knowledge of AI and wireless positioning technologies. For AI technologies, we summarize the basic principles, introduce classical neural network models, and discuss key AI algorithms. For wireless positioning, we review various application scenarios, emerging wireless technologies, and fundamental concepts, including channel models and basic methodologies.
  • 3GPP Standardization Progress: We analyze the evolution of cellular positioning within the 3rd Generation Partnership Project (3GPP) framework, summarizing its advancements from early implementations to current 5G standards. We also discuss the role of key performance indicators (KPIs) in wireless positioning and highlight the latest advancements in AI/ML-driven positioning solutions within the 3GPP standards.
  • SOTA Research of AI-Driven Positioning: Based on 3GPP-defined frameworks, we summarize the state-of-the-Art (SOTA) in AI/ML-assisted positioning methods and direct AI/ML positioning methods. For AI/ML-assisted positioning methods, we focus on SOTA techniques for positioning parameter estimation, including AI-based LOS/NLOS detection, TOA/Time-difference-of-arrival (TDOA) estimation, and angle estimation algorithms. For direct AI/ML positioning methods, we classify and summarize the methods into fingerprint-based positioning, knowledge-assisted AI positioning, and channel charting based positioning, introducing the latest progress in each.
  • Datasets for Cellular Positioning: We review publicly available datasets for cellular positioning, analyzing their scenarios and unique characteristics. These datasets are essential for benchmarking AI-driven wireless positioning systems and provide critical resources for researchers seeking to validate and improve their algorithms.
  • Challenges and Future Directions: We discuss the major challenges in AI-driven wireless positioning, including the scarcity of high-quality training data, the computational complexity of AI algorithms, and the model Generalization. Finally, we propose potential research directions and opportunities, providing guidance from both the perspectives of wireless technology development and AI technology advancements.
Therefore, the structure of this paper is illustrated in Fig. 1. In Sec. II, we present the fundamentals of AI technology, while Sec. III focuses on the fundamentals of wireless positioning. Sec. IV summarizes the progress in 3GPP AI/ML positioning standards. The state-of-the-art advancements are reviewed in Sec. V for AI/ML-assisted positioning methods and in Sec. VI for direct AI/ML positioning methods. In Sec. VII, we investigate publicly available datasets for wireless positioning, analyzing their characteristics and application scenarios. Sec. VIII identifies the challenges and opportunities in AI-driven wireless positioning. Finally, Sec. IX concludes this paper.
II Fundamentals of AI Technology
Advanced AI models and algorithms play a critical role in enhancing the accuracy and efficiency of AI-driven wireless positioning systems. In this section, we will provide a review of classic AI models and algorithms.
II-A Overview of AI Models
(a) FCNN architecture.
(b) CNN architecture.
(c) LSTM architecture.
(d) Transformer architecture.
Figure 2: Architectures of the FCNN, CNN, LSTM and Transformer.
In the realm of AI, neural network architectures play a pivotal role in enabling machines to learn from data and make predictions or classifications. This subsection introduces four mainstream neural network models commonly applied in AI-driven wireless positioning systems: fully-connected neural networks (FCNNs), convolutional neural networks (CNNs), long short-term memory networks (LSTMs), and Transformers. The architecture of these models are shown in Fig. 2.
II-A1 FCNNs
FCNNs are the basic building blocks of deep learning. In an FCNN, every neuron in one layer is connected to every neuron in the subsequent layer, creating a fully connected architecture [22, 23, 3]. The network structure of a Fully Connected Neural Network (FCNN) is illustrated in Fig. 2(a). In an FCNN layer with $m$ input neurons and $n$ output neurons, each output neuron is represented as a weighted sum of all neurons from the preceding layer.
FCNNs can be employed to tackle regression, prediction, and classification tasks. When disregarding computational performance constraints, configuring deeper network models often leads to better performance. However, being a linear neural network, FCNNs inevitably encounter issues with poor accuracy when dealing with nonlinear datasets. Consequently, subsequent efforts have often focused on augmenting the nonlinear characteristics of neural networks.
II-A2 CNNs
CNNs are a type of feedforward neural network featuring convolutional operations and deep structures. They are widely used in image processing [24, 25], speech recognition [26, 27], and natural language processing [28]. As shown in Fig. 2(b), in a CNN, the convolution operation applies a set of shared weights (convolution kernels) to an input feature map to produce an output feature map.
CNNs excel at extracting local features through their convolutional layers, leveraging translation invariance to ensure efficient and accurate feature representation. Compared to FCNNs, CNNs reduce the number of parameters by sharing weights across spatial locations, making them computationally efficient for high-dimensional data [29]. However, due to their focus on local regions, CNNs inherently struggle to model global relationships, particularly at a fine-grained level. This limitation has led to the integration of complementary techniques, such as attention mechanisms or global pooling layers, to enhance their ability to capture long-range dependencies.
II-A3 LSTMs
CNNs are effective for spatial modeling but are limited in capturing temporal dependencies. To address this, LSTM networks [30, 31], a type of recurrent neural network (RNN), are designed to model sequential data by introducing gate mechanisms to regulate information flow. The schematic diagram of an LSTM is shown in Fig. 2(c). An LSTM processes input sequences using mechanisms such as the forget gate to discard irrelevant past information, the input gate to incorporate new data, and the output gate to produce meaningful hidden states.
These features enable LSTMs to retain essential long-term dependencies while adapting to new input dynamically. However, their recurrent nature and additional parameters make them computationally intensive compared to simpler feedforward networks. Beyond LSTM, other neural network models like gated recurrent unit (GRU) [32], Bidirectional RNN (BiRNN) [33], temporal convolutional network (TCN) [34], and Transformers [35] have been extensively used for time series tasks. GRU offers efficiency with fewer parameters, while BiRNN captures bidirectional dependencies. TCN achieves scalable performance by replacing recurrence with causal convolutions. Transformers leverage attention mechanisms to model long-term dependencies and are increasingly popular for time series applications.
II-A4 Transformer
Transformers are groundbreaking neural network architectures built entirely on the attention mechanism, which have revolutionized AI across multiple domains [35]. Their unified encoder-decoder structure and multimodal capabilities make them highly versatile and adaptable. As shown in Fig. 2(d), the Transformer is mainly composed of $N_{e}$ cascaded encoders and $N_{d}$ cascaded decoders. The core of the Transformer is the self-attention mechanism, which assigns dynamic weights to different parts of the input sequence, enabling the model to focus on the most relevant features [36].
Transformers streamline feature processing with a unified architecture, but their $O(n^{2})$ computational complexity for attention mechanisms imposes high resource requirements, particularly for long sequences [37]. Moreover, the reliance of Transformers on large datasets for training can limit their efficiency on smaller datasets or simpler tasks, where models like CNNs may perform better.
II-B Overview of AI Algorithms
(a) Transfer Learning.
(b) Continual Learning.
(c) Few-Shot Learning.
(d) Generative Adversarial Network.
Figure 3: Schematic diagram of Transfer learning, Continuous learning, Few-shot learning algorithms, and Generative adversarial network.
In this subsection, we will introduce the AI algorithms commonly used in wireless AI positioning, including transfer learning, continual learning, few-shot learning, and generative adversarial networks (GANs). The schematic diagram of each algorithm is shown in Fig. 3.
II-B1 Transfer Learning
Transfer learning is an AI technique that utilizes knowledge from a source domain or task to enhance learning in a target domain or task [38, 39]. This method is effective when labeled data in the target domain is limited or costly to acquire, while sufficient data or pre-trained models are available in a related source domain. As shown in Fig. 3(a), transfer learning achieves this by reusing representations from pre-trained models and adapting them to new tasks through two main steps. First, in the pre-training phase, the pre-trained model acts as a feature extractor, capturing transferable patterns from the source task. During the fine-tuning phase, the model is further trained on the target domain data, aligning it with the specific requirements of the new task by adjusting part or all of its parameters. This process enables AI models to transfer knowledge from the source domain to the target domain, significantly improving performance, especially when labeled data in the target domain is limited [39].
Transfer learning has shown great promise in wireless positioning, where collecting labeled real-world data is often challenging. For instance, knowledge from simulated datasets can be transferred to real-world scenarios, significantly reducing the need for extensive labeled data while improving system performance [40, 41, 42]. Similarly, a model trained on data from one floor of a building can be adapted to other floors, thereby lowering deployment costs. Furthermore, transfer learning can address challenges in cross-band and cross-protocol deployments, enabling models to generalize across different frequency bands or wireless communication protocols. However, the effectiveness of transfer learning in wireless positioning is closely tied to the similarity between the source and target domains. A significant discrepancy between the two domains can reduce transfer learning’s effectiveness, while a small target dataset may lead to overfitting during fine-tuning.
II-B2 Continual Learning
Continual learning [43, 44], also known as lifelong learning [45] or incremental learning [46, 47], refers to the capability of AI models to continuously learn and accumulate knowledge from past tasks, facilitating adaptation to new and emerging tasks. The schematic diagram of the continual learning algorithm is shown in Fig. 3(b). This process mimics human learning by building and retaining an evolving body of knowledge, ensuring that previously learned information is not forgotten when encountering new tasks. While continual learning shares similarities with transfer learning in leveraging prior knowledge, its objective is distinct: continual learning emphasizes long-term knowledge accumulation, whereas transfer learning focuses on optimizing a specific target task without considering long-term retention.
Continual learning is essential in wireless positioning systems, which operate in dynamic environments and constantly evolving deployment scenarios. For instance, continual learning enables positioning models to adapt to new wireless network configurations or expanded coverage areas without losing effectiveness in previously trained conditions [48]. It also addresses challenges such as adapting to evolving wireless standards, new frequency bands, changing signal propagation conditions, and hardware upgrades. Despite its potential, continual learning faces challenges, including the risk of overfitting to new tasks when data is limited and the computational overhead associated with regularization or replay-based approaches. Developing efficient and scalable continual learning methods remains a critical area for advancing wireless positioning systems.
II-B3 Few-Shot Learning
As shown in Fig. 3(c), few-shot learning aims to train effective models using a limited number of samples [49, 50]. This approach reduces the dependency on large datasets, significantly lowering the cost of data annotation and collection while maintaining competitive model performance. In scenarios where acquiring sufficient labeled data is challenging, Few-shot learning offers a promising solution by leveraging innovative learning paradigms to extract meaningful insights from sparse data. Few-shot learning models can be broadly categorized into three types: model-based, metric-based, and optimization-based approaches [49, 50].
Few-shot learning is also important for wireless positioning systems, where labeled data is often scarce, and data collection is costly. For example, it can train models for positioning in new environments or deployment configurations using minimal labeled data. However, several challenges remain. The success of Few-shot learning depends heavily on the quality and representativeness of the support set used during training. In highly diverse or complex wireless environments, identifying representative samples can be difficult. Moreover, the limited number of samples can make Few-shot learning susceptible to overfitting, especially in noisy or heterogeneous datasets. Addressing these challenges requires careful design of learning paradigms and effective strategies for generalization and robustness in wireless positioning applications.
II-B4 Generative Adversarial Networks
As shown in Fig. 3(d), GANs comprise two primary components: a Generator and a Discriminator. These two components engage in a competitive learning process where the Generator strives to produce data that closely resembles real data, while the Discriminator aims to distinguish between real and generated data. This adversarial training framework enables both models to iteratively improve their performance [51, 52].
In wireless systems, GANs are a powerful tool for data augmentation and system optimization. Their capability to generate synthetic data resembling real-world conditions is particularly beneficial in scenarios where collecting real data is expensive or logistically challenging. For example, GANs can synthesize wireless channel data under varying environmental conditions, enabling the training of robust AI models for positioning without requiring extensive real-world data collection. Moreover, GANs can assist in addressing class imbalance issues and enhancing the diversity of training datasets, thereby improving model generalization and reliability.
II-B5 Other Advanced AI Techniques
Beyond the AI algorithms discussed, advanced techniques such as knowledge distillation [53], ensemble learning [54], and self-supervised learning [55] offer additional avenues to enhance wireless positioning systems. For example, knowledge distillation involves transferring knowledge from a large, complex model (teacher) to a smaller, more efficient model (student), improving inference speed and scalability while maintaining high accuracy. By combining predictions from multiple models, ensemble learning improves robustness and accuracy, particularly in heterogeneous environments. Self-supervised learning leverages unlabeled data by generating pseudo-labels through pretext tasks, reducing the reliance on labeled datasets and enabling models to learn effective feature representations.
As AI technologies continue to evolve, these advanced techniques and future innovations will play a critical role in overcoming challenges and achieving higher accuracy, efficiency, and robustness in wireless positioning systems. Integrating these methodologies will remain an ongoing focus in research and practical deployments.
III Fundamentals of Wireless Positioning
Figure 4: AI-driven wireless positioning technologies and applications.
In this section, we first discuss the application scenarios of wireless positioning, and then list emerging wireless technologies. Finally, we introduce the basics of wireless positioning in detail, including channel models and algorithm foundations. AI-driven wireless positioning technology and applications are shown in Fig. 4.
III-A Applications of Wireless Positioning
Wireless positioning has emerged as a transformative technology, enabling a wide range of applications that improve efficiency, safety, and user experience across various domains. From intelligent transportation systems to healthcare and beyond, the ability to accurately determine the location of devices and objects has unlocked significant advancements in modern technologies. In this subsection, we will introduce the importance of wireless positioning from several representative application scenarios.
III-A1 Intelligent Transportation
Wireless AI positioning is pivotal in intelligent transportation systems, enabling precise vehicle localization that optimizes traffic flow, reduces congestion, and enhances road safety [56, 57, 58]. Wireless positioning facilitates fleet management and route optimization for public transportation and ride-sharing services. Accurate location data also enhances road safety by enabling real-time hazard detection and proactive alerts to drivers [59].
III-A2 Autonomous Driving
The success of autonomous driving depends heavily on high-precision positioning, which ensures vehicles can navigate safely and efficiently in complex environments. Wireless AI positioning offers the accuracy needed for lane-level navigation, seamless obstacle avoidance, and real-time decision-making in dynamic traffic scenarios [60, 61, 62]. Furthermore, positioning data is integral to vehicle-to-everything (V2X) communication, supporting synchronized interactions between vehicles and infrastructure to enhance traffic efficiency and safety.
III-A3 Extended Reality
In XR applications, precise positioning of devices and users is crucial for creating seamless and immersive experiences. Whether in augmented reality (AR) navigation [63], virtual reality (VR) training simulations [64, 65, 66], or multiplayer gaming environments [67], wireless positioning ensures accurate tracking of movements and spatial relationships. This enhances the realism of interactive simulations and allows for seamless transitions between indoor and outdoor environments in XR-based applications.
III-A4 Indoor Tracking and Navigation
Wireless positioning systems transform navigation within complex indoor environments [68, 69, 70, 71], such as airports, shopping malls, and hospitals. These systems provide users with precise guidance to specific locations, improving efficiency and saving time in environments where traditional GPS signals are unreliable. For businesses, indoor navigation enhances operational workflows and provides opportunities for personalized user engagement, while improving overall customer experience.
III-A5 Public Safety
Public safety applications greatly benefit from wireless AI positioning, particularly in emergency scenarios where speed and accuracy are critical [72, 73]. By locating individuals in disaster-stricken or high-risk areas, such as collapsed buildings or burning structures, positioning systems streamline rescue operations and improve the success rates of life-saving interventions [74].
III-A6 Internet of Things
Wireless positioning underpins the functionality of internet of things (IoT) ecosystems, enabling efficient management and monitoring of smart devices across various environments [75, 19]. In smart homes, factories, and agricultural settings, precise location data improves device interconnectivity and automation. By tracking equipment, inventory, and resources in real time, IoT-enabled positioning reduces operational inefficiencies and supports smarter decision-making.
III-A7 Security and Surveillance
Security and surveillance systems utilize wireless AI positioning to monitor the real-time locations of personnel and assets in sensitive areas such as prisons, factories, and warehouses. By integrating real-time positional data, these systems enhance safety protocols and operational efficiency, ensuring the secure management of critical environments [76, 77].
III-A8 Sports and Motion Analysis
In sports and health domains, wireless positioning facilitates detailed motion analysis by tracking athletes’ positions and trajectories [78]. This data supports the optimization of training programs and enhances performance evaluations, providing athletes and coaches with actionable insights to refine techniques and strategies.
III-A9 Healthcare
In healthcare settings, wireless positioning enables real-time tracking of patients, staff, and equipment, improving the delivery of medical services and overall operational efficiency [79]. For instance, real-time location monitoring in critical care units, such as intensive care units or operating rooms, allows for rapid response to emergencies, potentially saving lives in time-sensitive situations.
III-A10 UAV positioning
UAVs depend on wireless AI positioning for accurate navigation, collision avoidance, and stable operation in applications like surveillance, delivery, and disaster response [80, 81, 82]. In GPS-denied environments, such as indoors or dense urban areas, wireless positioning systems using technologies like UWB and 5G enable precise localization. AI-driven approaches further enhance positioning accuracy by integrating wireless signals with inertial and vision-based data, supporting tasks like agricultural spraying and infrastructure inspection. Despite challenges like interference and dynamic environments, advancements in multi-sensor fusion and cooperative positioning continue to improve UAV efficiency and safety.
III-B Emerging Wireless Technologies
Emerging wireless technologies are reshaping the landscape of wireless positioning by introducing novel architectures and signal propagation characteristics that enhance accuracy, scalability, and reliability. Key advancements include cell-free networks, massive multiple-input multiple-output (MIMO), non-terrestrial networks (NTNs), THz communications, and RIS. Subsequently, we will introduce several key technologies.
III-B1 Massive MIMO
Massive MIMO technology, with its deployment of large-scale antenna arrays at base stations (BSs), significantly improves spatial resolution and beamforming precision. This enhances the ability to estimate parameters such as the angles-of-arrival (AOA), direction-of-arrival (DOA), angle-of-departure (AOD), and TOA. The high spatial diversity provided by massive MIMO also mitigates the effects of multipath propagation, making it a cornerstone for precise positioning in 5G and beyond. Additionally, Flexible Location MIMO [83], based on fluid antennas and movable antennas, represents a novel antenna paradigm. These antennas can move dynamically to avoid interference and obstacles, ensuring fast and reliable wireless transmission.
III-B2 Cell-Free
Cell-free networks utilize a distributed architecture where a dense network of access points (APs) collaborates to serve users [84]. This architecture mitigates inter-cell interference and provides seamless coverage, enhancing positioning accuracy, especially in dense urban or indoor environments. By exploiting the joint processing of signals from multiple APs, cell-free networks enable highly precise multi-AP positioning [84]. Additionally, the system’s inherent flexibility supports adaptive coverage and robust performance in dynamic scenarios.
III-B3 Non-Terrestrial Networks
NTNs, encompassing satellite constellations, high-altitude platforms, and UAVs, extend wireless connectivity to remote and underserved areas [85]. These networks complement terrestrial systems by providing wide-area coverage and redundancy. For positioning, NTNs offer unique advantages, such as a clear LOS and large coverage areas, enabling applications in maritime, aviation, and disaster recovery scenarios.
III-B4 Terahertz Networks
THz communication operates in the spectrum above 100 GHz, offering ultra-wide bandwidths that allow high-resolution parameter estimation. The short wavelength of THz signals enables precise localization through fine-grained TOA and AOA measurements. These networks are particularly suited for ultra-dense environments and applications requiring sub-centimeter positioning accuracy, such as industrial automation and XR [18].
III-B5 reconfigurable intelligent surfaces
RIS technology uses programmable metasurfaces to manipulate electromagnetic waves, creating controllable propagation environments. By dynamically adjusting reflection, refraction, or scattering, RIS can enhance signal quality, and improve positioning accuracy in challenging environments. For example, RIS can redirect signals to establish virtual LOS paths or amplify weak signals in dense urban areas [86]. The integration of RIS with AI-driven positioning systems promises unprecedented levels of control and precision, though challenges such as hardware scalability and real-time optimization remain.
III-B6 Near-field communication
Near-field Communication enhances wireless positioning by leveraging spherical wavefronts in close proximity environments, enabling precise parameter estimation like phase and amplitude variations [87]. This improves accuracy in technologies such as massive MIMO and RIS, making it ideal for applications like indoor navigation and industrial automation. However, near-field communication increases computational complexity and requires advanced models to capture wavefront characteristics, presenting both opportunities and challenges for next-generation wireless networks.
The integration of these technologies significantly enhances wireless positioning capabilities. By leveraging their unique characteristics, such as improved spatial resolution in massive MIMO, flexibility in RIS, and wide coverage in NTNs, these advancements address the limitations of conventional systems. Furthermore, combining these technologies with AI holds promise for achieving robust, scalable, and ultra-precise positioning in next-generation wireless networks.
III-C Basics of Wireless Positioning
Wireless positioning technology fundamentally involves the estimation of a device or object’s location through the measurement of the propagation characteristics of wireless signals between a transmitter and a receiver. Path loss is a key and intuitive factor in this process, which describes the attenuation of signal power over distance [88]. Path loss can be expressed using models like the free-space path loss equation:
$PL(d)=PL(d_{0})+10n\log_{10}\left(\frac{d}{d_{0}}\right),$ (1)
where $PL(d)$ represents the path loss as a function of distance $d$, $PL(d_{0})$ is the path loss at a reference distance $d_{0}$, and $n$ is the path loss exponent, which depends on the environment.
Based on path loss, signal strength inherently reflects the spatial relationship between the transmitter and receiver. This characteristic makes received signal strength (RSS) metrics, such as received signal strength indicator (RSSI) and reference signal received power (RSRP), valuable measurements for wireless positioning. RSSI represents a measure of the total power present in a received radio signal, encompassing both the desired signal and any interference or noise [89]. It is widely used in wireless technologies such as WiFi, Zigbee, and Bluetooth. On the other hand, RSRP is a metric specific to Long-term Evolution (LTE) and 5G networks [90]. It measures the power of the reference signal transmitted by BSs, providing a more precise and isolated indicator of signal quality and strength for cellular systems. These metrics allow wireless systems to infer distances between devices and BSs, forming the foundation for various positioning techniques. By relating measured RSSI or RSRP to known transmission power levels via path loss models, the distance between the transmitter and receiver can be estimated. These distance estimates form the foundation of range-based positioning techniques, such as trilateration [91], which will be discussed later.
However, in the practical system, the use of RSSI and RSRP is affected by factors such as multipath propagation, environmental obstacles and noise, which may introduce errors into the distance estimation [92]. To better characterize the wireless transmission environment, a more detailed metric, channel state information (CSI), is often employed [93]. CSI provides a comprehensive characterization of the wireless channel response between the transmitter and receiver, including both amplitude and phase information for each subcarrier in the frequency domain. Unlike RSSI and RSRP, which are scalar measures, CSI is a multidimensional metric that captures the channel’s response across frequencies or subcarriers. Additionally, the detailed frequency-domain information provided by CSI allows for advanced modeling techniques, such as ML-based approaches, to further enhance positioning accuracy in dynamic and challenging environments [94].
Figure 5: MIMO-OFDM-based wireless positioning systems.
As an illustrative example, we consider a MIMO-orthogonal frequency-division multiplexing (OFDM) system to model CSI in a wireless positioning system. As shown in Fig. 5, in the MIMO-OFDM system, the BS is equipped with a uniform plane array, comprising $N_{a}$ antennas with $N_{a}^{r}$ in each row and $N_{a}^{c}$ in each column. The user equipment (UE) is equipped with a single omnidirectional antenna. The system contains $N_{s}$ subcarriers, each with a bandwidth of $B_{s}$. The locations of the BS and the UE can be represented as $\bm{p}^{ue}$ and $\bm{p}^{bs}$ respectively. The channel frequency response between the BS and the UE for subcarrier $s$ can be written as [95]
$\mathbf{H}_{s}=\sum_{l=1}^{L}\alpha_{l}\mathbf{a}(\theta_{l},\varphi_{l})e^{-j2\pi f_{s}\tau_{l}},$ (2)
where $f_{s}$ is the subcarrier frequency, $\alpha_{l}$ and $\tau_{l}$ denote the complex gain and delay of path $l$, and $\mathbf{a}(\theta_{l},\varphi_{l})$ is the two-dimensional array response matrix. For the LOS path, $\tau_{1}$ and $(\theta_{1},\varphi_{1})$ represent the delay and AOA, respectively.
Specifically, $\mathbf{a}(\theta_{l},\varphi_{l})$ can be expressed as
$\mathbf{a}(\theta_{l},\varphi_{l})=\mathbf{a}_{h}(\theta_{l},\varphi_{l})\otimes\mathbf{a}_{v}(\varphi_{l}),$ (3)
where $\mathbf{a}_{h}(\theta_{l},\varphi_{l})$ and $\mathbf{a}_{v}(\varphi_{l})$ represent the elevation and azimuth components, respectively, and are defined as
$\mathbf{a}_{h}(\theta_{l},\varphi_{l})$ $=$ $\left[1,e^{-j2\pi\frac{d^{h}}{\lambda}\sin\varphi_{l}\cos\theta_{l}},\ldots,\right.$ (4)
$\quad\left.e^{-j2\pi(M-1)\frac{d^{h}}{\lambda}\sin\varphi_{l}\cos\theta_{l}}\right]^{\mathrm{T}},$
$\mathbf{a}_{v}(\varphi_{l})$ $=$ $\left[1,e^{-j2\pi\frac{d^{v}}{\lambda}\cos\varphi_{l}},\ldots,\right.$ (5)
$\quad\left.e^{-j2\pi(N-1)\frac{d^{v}}{\lambda}\cos\varphi_{l}}\right]^{\mathrm{T}}.$
Here, $\lambda$ is the wavelength, and $d^{h}$ and $d^{v}$ are the inter-antenna spacings in the horizontal and vertical directions, respectively, where $d^{h}=d^{v}=\frac{\lambda}{2}$. The full channel matrix in the frequency domain can then be expressed as $\mathbf{H}\triangleq[\mathbf{H}_{1},\cdots,\mathbf{H}_{s},\cdots,\mathbf{H}_{N_{s}}]$.
The channel frequency response (CFR) matrix $\mathbf{H}$ contains rich spatial and temporal information that directly relates to the physical characteristics of the propagation environment. Therefore, the research goal of wireless positioning is to determine a function $\mathcal{F}(\cdot)$ that can estimate an accurate position of the device based on the channel response $\mathbf{H}$. Therefore, the wireless positioning problem can be expressed as the following optimization problem:
$\min_{\mathcal{F(\cdot)}}$ $\quad||\hat{\bm{p}}^{ue}-\bm{p}^{ue}||_{2}^{2},$ (6)
s.t. $\quad\hat{\bm{p}}^{ue}=\mathcal{F}(\mathbf{H}),$ (7)
where $\bm{p}^{ue}$ and $\hat{\bm{p}}^{ue}$ represent the estimated position and the true position, respectively. However, this is a challenging problem due to the complexity of the wireless propagation environment, which includes factors such as multipath propagation, noise, and NLOS conditions.
To find an optimal $\mathcal{F}(\cdot)$, existing research generally adopts two main approaches. The first approach relies on the geometric relationships between the transmitter and the device, focusing on extracting geometric features like TOA and AOA from the CSI. These features are then utilized to calculate the UE’s position based on geometric principles.
The second approach uses artificial intelligence algorithms for direct end-to-end localization without explicitly estimating intermediate geometric parameters. For instance, in fingerprint-based positioning, raw CSI data or its derived features are used as inputs to AI models, which then output the position. These data-driven methods learn complex mapping functions between CSI and location, enabling high-precision positioning without relying on geometric computations.
In the following sections, we provide a detailed introduction to the principles of these methods, including their advantages, limitations, and typical application scenarios.
(a) TOA-based positioning.
(b) TDOA-based positioning.
(c) Angle-based positioning.
(d) Fingerprint-based positioning.
Figure 6: Schematic diagram of positioning technology based on TOA, TDOA, angle and fingerprint.
III-C1 Ranging-based Positioning
High-precision positioning can be achieved by estimating the distances between a device and multiple BSs. With geometric relationships, accurate positioning is possible when distance estimations to at least three BSs are available. As shown in Fig. 6(a), the intersection of circles representing these distances identifies the device’s location. Adding more BSs and improving distance estimation accuracy further enhance positioning precision.
As previously discussed, one way to measure the distance between the BSs and the device is by fitting the relationship between the RSS and the distance, thus allowing distance estimation using RSS [96, 97, 98, 99, 100]. The same principles apply to methods based on RSRP in cellular systems and RSSI in WiFi systems. However, the propagation characteristics of wireless signals vary across different frequency bands, and the wireless propagation environment fluctuates over time. These factors make it challenging to determine an accurate path loss function. Factors such as multipath propagation, interference, and NLOS conditions further degrade the accuracy of RSS-based methods. Consequently, although RSS-based positioning is simple and cost-effective, it is not easy to achieve high-precision positioning.
To overcome the limitations of RSS-based methods, TOA estimation provides a more precise approach. TOA, also referred to as time-of-flight (TOF) [14], is used to measure the time taken for a wireless signal to propagate from the transmitter to the receiver. Based on the TOA measurement, the distance $d$ can then be calculated as $d=c\cdot t_{toa}$, where $t_{toa}$ is the signal propagation time, equal to the propagation delay of the LOS path $\tau_{1}$, and $c\approx 3\times 10^{8}$ m/s is the speed of light.
In practice, to estimate the TOA, BS transmits a predefined pilot signal, which is received by the device. The device performs channel estimation using the received signal and then estimates the TOA from the estimated channel [101, 102, 103, 104]. The TOA estimation process can be expressed as a function $\mathcal{F}_{toa}(\cdot)$ that maps the estimated channel $\hat{\mathbf{H}}$ to the TOA:
$t_{toa}=\mathcal{F}_{toa}(\hat{\mathbf{H}}).$ (8)
$\hat{\mathbf{H}}$ represents the estimated CSI, and $\mathcal{F}_{toa}(\cdot)$ denotes signal processing algorithms used to extract the TOA from the channel response. The accuracy of TOA-based distance estimation is influenced by several factors, including the pilot signal bandwidth, signal-to-noise ratio (SNR), and the sampling rate of the receiver. The Cramer-Rao lower bound (CRLB) can be used to describe the variance of the TOA estimation error as a function of the bandwidth $B$ and SNR [104, 105, 106, 107]:
$\text{Var}(t_{toa})\propto\frac{1}{B^{2}\cdot\text{SNR}}.$ (9)
This indicates that increasing the bandwidth and SNR enhances the accuracy of TOA estimation, albeit at the cost of requiring more advanced hardware and signal processing.
To improve TOA estimation accuracy, advanced algorithms like multiple signal classification (MUSIC) [108] and estimation of signal parameters via rotational invariance techniques (ESPRIT) [109] are widely used. MUSIC identifies peaks in the signal’s spatial spectrum, while ESPRIT leverages the rotational invariance of the signal subspace for precise estimation. However, these algorithms still face the challenge of robustness in different environments. Additionally, TOA-based positioning requires precise clock synchronization between the BS and UEs [110, 111, 112, 113], as clock misalignment introduces an unknown offset in TOA measurements. This offset introduces a systematic error in the distance estimation. Therefore, the measured TOA can be expressed as
$t_{toa}^{\text{measured}}=t_{toa}+\Delta t_{clk}+e_{toa},$ (10)
where $\Delta t_{clk}$ and $e_{toa}$ represents the clock synchronization error and estimation error.
To mitigate the effects of clock synchronization errors, alternative methods such as round-trip time (RTT) [114, 115, 116] estimation are often employed. RTT measures the total round-trip delay of the signal, effectively eliminating the need for synchronization between the BS and the device. Since the RTT is calculated using the clock in the transmitter, it effectively cancels out the impact of clock misalignment.
Moreover, when BSs are synchronized with each other, TDOA-based positioning technique [117, 118, 119] provides an effective alternative to address the lack of synchronization between the BS and the device. TDOA measures the difference in arrival times of a signal at multiple BSs. For example, the TDOA can be expressed as
$\Delta t_{tdoa}=$ $\ t_{toa,i}-t_{toa,j}$
$=$ $\ t_{toa,i}+\Delta t_{clk,i}+e_{toa,i}-t_{toa,j}+\Delta t_{clk,j}+e_{toa,j}$
$=$ $\ t_{toa,i}-t_{toa,j}+e_{tdoa}^{*},$ (11)
where $t_{toa,i}$ and $t_{toa,j}$ are the TOAs from the device to BS $i$ and BS $j$, respectively, and $e_{tdoa}^{*}$ represents the combined error. In this case, the error caused by the desynchronization can be eliminated. Similar to the TOA-baed positioning, TDOA-based positioning, as shown in Fig. 6(b), can draw a hyperbola based on the TDOA between each pair of BSs and the device. When the number of BSs exceeds three, the region where the device is located can be determined. Similar to TOA-based positioning techniques, the more BSs involved in the positioning process and the more accurate the TDOA estimation, the higher the positioning accuracy of the device. In cellular positioning systems (LTE and 5G), TDOA-based techniques are also referred to as observed time difference of arrival (OTDOA) [104, 105, 120].
III-C2 Angle-based positioning
Angle-based positioning techniques estimate the location of a device by determining angular parameters, specifically the AOA [121, 122, 123] and AOD [124, 125, 126], using the spatial characteristics of signals received or transmitted by antenna arrays.
The AOA indicates the direction from which a signal is received at the receiver, typically measured using uplink signals, while the AOD denotes the direction in which a signal is transmitted from the source, generally derived from downlink channels. For AOA estimation, For 3D positioning using AOA, it is crucial to estimate the 3D angular parameters, including the elevation angle $\theta_{1}$ and azimuth angle $\phi_{1}$ of the LOS path. Based on the channel model provided above, the AOA measurement error can be expressed as
$[\theta_{1},\phi_{1}]=\mathcal{F}_{aoa}(\hat{\mathbf{H}}).$ (12)
For 2D positioning, we only need to estimate $\theta_{1}$. The CRLB for the 2D angle estimates can be expressed as [107]
$\text{Var}(\theta_{1})\propto\frac{1}{(2\pi d_{a}/\lambda)^{2}\text{SNR}\sin^{2}{\theta_{1}}N(N-1)(2N-1)},$ (13)
where $d_{a}$ is the inter-antenna spacing, $N$ is the number of antennas in the array. The estimation method of AOD is similar to that of AOA.
Similar to ranging-based positioning techniques, as illustrated in Fig. 6(c), angle-based positioning estimates the position of a device by using the spatial direction between the device and multiple BSs. A straightforward approach involves beam scanning, where the beam direction with the maximum received power is selected as the AOA estimation result. In addition, advanced algorithms, such as MUSIC [127, 128] and ESPRIT [129, 130], are widely employed. These methods leverage the spatial characteristics of the received signal, enabling angular resolution that exceeds the physical limitations of the antenna array.
The primary advantage of angle-based positioning is its independence from time synchronization between the BSs and the device, which is a crucial requirement for time-based methods like TOA and TDOA. However, the performance of angle-based positioning greatly depends on the angular resolution and precision of the antenna arrays, as well as the algorithms used for AOA and AOD estimation. Higher resolution arrays and advanced signal processing techniques can significantly improve positioning accuracy but also demand more advanced hardware and increased computational resources.
III-C3 Fingerprint-based Positioning
Fingerprint-based positioning establishes a mapping between wireless network measurements and spatial locations by constructing an offline database or training AI models. It involves deriving a mapping function $\mathcal{F}{fp}(\cdot)$ from an offline dataset $D{train}={\bm{x}_{i},\bm{p}_{i}}$, where $\bm{x}_{i}$ represents wireless fingerprint features and $\bm{p}i$ denotes the corresponding UE position. The function $\mathcal{F}{fp}(\cdot)$ is obtained through database creation or AI model training.
As depicted in Fig. 6(d), fingerprint-based positioning typically consists of two main phases: the offline phase and the online phase. During the offline phase, a detailed survey of the environment is performed to collect wireless signal parameters, which are then used to construct a wireless fingerprint database, also known as a radio map. The key steps involved are as follows:
  • Offline Data Collection: Wireless signal characteristics, such as RSSI [131, 132, 133], RSRP [134, 135, 136, 137, 138, 139], CSI [140, 141, 142, 143, 144], or TOA measurements [145, 146], are collected at predefined reference points (RPs) within the target area. These RPs are typically arranged in a grid or sampled randomly.
  • Fingerprint Feature Extraction: The collected raw data is processed to extract meaningful features, such as average signal strength or variance. Alternatively, the raw signal data itself can also be used directly as a fingerprint.
  • Fingerprint Database Creation: The extracted features for each RP are compiled into a fingerprint database, denoted as $D_{train}$. Each database entry comprises a pair, $\bm{x}_{i}$ (signal features) and $\bm{p}i$ (corresponding RP coordinates), forming a virtual map of the fingerprint. Beyond traditional database construction [147], AI/ML models can be utilized to learn a mapping function $\mathcal{F}{fp}(\cdot)$ that directly correlates input features to output positions, serving as an alternative to conventional databases.
In the online phase, the goal is to determine the position of an unknown device by comparing its real-time signal measurements with the pre-established fingerprint database. Therefore, For a UE with fingerprint features $\bm{x}$, its estimated position can be expressed as
$\hat{\bm{p}}=\mathcal{F}_{pf}(\bm{x}).$ (14)
This position estimation process includes:
  • Real-Time Signal Measurement: The device requiring positioning continually collects wireless data at its current position, extracting the wireless fingerprint features.
  • Feature Matching and Position Estimation: Various matching algorithms are used to compare current signal measurements against entries in the fingerprint database. Location is estimated by identifying the RP whose fingerprints best match the observed signals. If an AI/ML model serves as the fingerprint mapping, the process involves inputting the extracted wireless features into the AI/ML model to output the estimated position directly.
  • Position Result Output: The result of the matching or model output is a predicted position, typically expressed as coordinates or descriptive position labels.
Fingerprint-based positioning relies on learning from a historical database of wireless measurements but involves high deployment costs due to extensive data collection and maintenance. Adapting AI models to map signal features to locations demands significant computational resources and expertise. Additionally, dynamic wireless environments can degrade positioning accuracy over time due to data aging and environmental changes. This necessitates periodic updates to the fingerprint database or retraining of the AI model to maintain high accuracy levels.
III-C4 Channel Charting based Positioning
Channel charting is a novel approach leveraging CSI to enable the pseudo-positioning of users by modeling relationships between different channels [148]. It learns a mapping from CSI to a lower-dimensional channel chart, where distances in the latent space represent dissimilarity metrics between channels [149, 150, 151, 152]. These distances reflect the proximity of devices in the actual propagation environment, with higher channel similarity corresponding to closer proximity on the channel chart [148]. Besides positioning, channel charting based techniques have been explored in various applications, such as SNR prediction [153], pilot assignment [154, 155], pilot reuse [156], beam tracking [157] and wireless resource optimization [158].
In contrast to fingerprint-based methods, which depend on pre-collected wireless fingerprints and external RPsto achieve absolute positioning, channel charting employs self-supervised learning to reduce the dimensionality of CSI data. Its objective is to uncover latent spatial relationships while minimizing the need for labeled data.
Channel charting, a form of manifold learning, encompasses classical methods such as multidimensional scaling (MDS), Sammon mapping [159], and t-SNE [160], alongside deep learning approaches like Siamese Neural Networks and triplet-based dimensionality reduction.
In Siamese-based channel charting systems, dissimilarity metrics define relationships between CSI samples in the latent space. The channel charting learning process aims to ensure that latent space distances preserve the high-dimensional relationships captured by the dissimilarity metrics. A commonly used MDS-based loss function can be given by [150]
$\mathcal{L}_{MDS}=\sum_{i,j}w_{ij}\left(\|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}-d(\mathbf{H}_{i},\mathbf{H}_{j})\right)^{2},$ (15)
where $\mathbf{z}_{i}$ and $\mathbf{z}_{j}$ are both the outputs of the neural network and the low-dimensional representations of the CSI samples $\mathbf{H}_{i}$ and $\mathbf{H}_{j}$. $d(\mathbf{H}_{i},\mathbf{H}_{j})$ is the dissimilarity metric in the original high-dimensional CSI space. $w_{ij}$ is a weight factor, often inversely proportional to $d(\mathbf{H}_{i},\mathbf{H}_{j})$, to prioritize preserving relationships for closer samples. This loss ensures that the distances in the latent space ($\|\mathbf{z}_{i}-\mathbf{z}_{j}\|$) align with the dissimilarity metrics derived from the high-dimensional CSI space, preserving spatial relationships. Two common measures of difference include:
  1. 1.
    Euclidean Distance [161, 162, 163]:
    $d_{E}(\mathbf{H}_{i},\mathbf{H}_{j})=\|\mathbf{H}_{i}-\mathbf{H}_{j}\|_{2},$ (16)
  2. 2.
    Cosine Similarity[164, 165, 166]:
    $d_{C}(\mathbf{H}_{i},\mathbf{H}_{j})=1-\frac{\langle\mathbf{H}_{i},\mathbf{H}_{j}\rangle}{\|\mathbf{H}_{i}\|_{2}\cdot\|\mathbf{H}_{j}\|_{2}},$ (17)
Triplet-based dimensionality reduction is another approach [164, 167, 165, 168]. In Triplet-based channel charting, relationships between CSI samples are learned through triplets $(\mathbf{H}_{i},\mathbf{H}_{j},\mathbf{H}_{k})$. Each triplet consists of an anchor ($\mathbf{H}_{i}$), a positive sample ($\mathbf{H}_{j}$) that is similar to the anchor, and a negative sample ($\mathbf{H}_{k}$) that is dissimilar. The objective is to map these samples into a latent space:
$\|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}<\|\mathbf{z}_{i}-\mathbf{z}_{k}\|_{2},$ (18)
where $(\mathbf{z}_{i},\mathbf{z}_{j},\mathbf{z}_{k})$ are the low-dimensional representations of $\mathbf{H}_{i},\mathbf{H}_{j},\mathbf{H}_{k}$, respectively.
The corresponding triplet loss function can be written as [167]
$\mathcal{L}_{\text{triplet}}=\sum_{i,j,k}\max\left(0,\|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}^{2}-\|\mathbf{z}_{i}-\mathbf{z}_{k}\|_{2}^{2}+\delta\right),$ (19)
where $\delta$ is a margin that enforces a minimum separation between positive and negative pairs.
As shown in Fig. 7, the left figure illustrates the physical space distribution of multiple sampling channels. In contrast, the right figure demonstrates the distribution of these points in the latent space after applying the channel charting approach. The relative proximity in the latent space correlates with the physical proximity of the devices.
(a) The actual physical space distribution of sampling points.
(b) Latent space distribution of sampling points.
Figure 7: Schematic diagram of channel charting. The relative proximity in the latent space correlates with the physical proximity of the devices.
The primary advantage of channel charting based positioning is its ability to avoid the use of external RPs, significantly reducing costs and complexity. Furthermore, channel charting continuously benefits from newly acquired channel information rather than being limited by static historical databases [169]. However, channel charting based positioning techniques rely heavily on the quality of the low-dimensional mapping, and there is often no universal metric to evaluate the effectiveness of these mapping algorithms [148]. Additionally, considering the highly nonlinear relationship between channel-based Dissimilarity metrics and physical distances, designing good metrics to improve positioning accuracy remains a significant challenge.
IV Progress in AI Positioning Standards
IV-A Evolution of 3GPP Standards for Positioning
1G networks were based on analog communication and lacked standardized positioning functionalities. Positioning was primarily achieved through signal strength measurements, such as RSS, which were used for optimizing BS selection, channel allocation, and handovers. Proprietary solutions, such as TDOA and AOA in advanced mobile phone system (AMPS) signals, supported emergency services and intelligent transportation systems [1, 170]. However, the overall positioning accuracy remained low.
2G networks introduced digital communications, and positioning technologies were initially standardized in the global system for mobile communications (GSM) [1], including cell-ID, time advance, enhanced observed time difference (E-OTD), and assisted GPS (A-GPS) [171]. These technologies significantly improved positioning accuracy, supported emergency services (such as E911), and laid the foundation for the development of subsequent positioning technologies.
Positioning technologies advanced further in 3G networks with the adoption of universal mobile telecommunications system (UMTS) and CDMA2000 standards. UMTS introduced techniques such as OTDOA, uplink time difference of arrival (UTDOA), and A-GPS [172]. radio frequency pattern matching (RFPM) was also employed to optimize performance [170]. CDMA2000 combined advanced forward link trilateration (AFLT) with A-GPS, leveraging network synchronization for efficient positioning [170]. These methods improved positioning accuracy to a range of 25–200 meters, addressing higher performance requirements in various applications.
With the introduction of LTE Release 9, positioning technologies were standardized for the first time. This standardization not only enhanced positioning accuracy but also addressed diverse scenarios and requirements. Positioning became an essential module for ensuring service quality, optimizing networks, and supporting emergency response. Key positioning methods introduced in LTE include [173]:
  • enhanced Cell-ID (E-CID): Improved accuracy by utilizing RSRP, and transmission time differences between mobile terminals and BSs.
  • OTDOA: Introduced positioning reference signal (PRS), transmitted in low-interference subframes to enable terminals to measure time differences and compute reference signal time difference (RSTD).
  • Assisted GNSS (A-GNSS): Supported GPS and other satellite navigation systems, with network assistance to provide differential corrections in signal-restricted environments.
To support cellular positioning, PRS signals were transmitted in dedicated subframes to minimize interference and enhance OTDOA performance. Furthermore, LTE positioning protocols, such as LTE positioning protocol (LPP), were defined to facilitate efficient communication between mobile terminals and positioning servers.
Subsequently, LTE-advanced (LTE-A) introduced UTDOA, RFPM [174], and enhanced OTDOA methods, improving positioning accuracy through multi-antenna transmission, heterogeneous networks and cross-network measurement [170]. The performance of PRS signals and cell hearability was also optimized. LTE-A also further enhanced positioning capabilities, introducing D2D-assisted positioning, MIMO vertical positioning and Bluetooth/WLAN/barometer positioning to meet positioning needs in multiple scenarios[175].
With the evolution of 5G technologies, positioning capabilities in 3GPP standards have advanced significantly to meet the growing demands of diverse applications. In Release 16, the groundwork for 5G positioning was laid by defining positioning use cases and service requirements [176, 177]. These efforts were aimed at addressing the initial needs for positioning in 5G networks. Subsequent studies focused on enabling new radio (NR) positioning techniques across both FR1 (sub-6 GHz) and FR2 (mmWave) bands. Key aspects included the specification of NR Downlink and Uplink reference signals to support techniques such as Downlink TDOA (DL-TDOA), Downlink AOD (DL-AOD), Uplink TDOA (UL-TDOA), Uplink AOA (UL-AOA), Multi-cell RTT, and E-CID. Additionally, the NR positioning protocol A (NRPPa) was introduced [178], providing a standardized communication framework for positioning in 5G networks.
To meet the demands of emerging 5G applications and vertical industries requiring higher accuracy, lower latency, and improved reliability, 3GPP launched the "Study on NR Positioning Enhancements" in Release 17 [179]. This study focused on solutions for high-accuracy, low-latency positioning that would also improve network and device efficiency. Key advancements included exploring positioning in in-coverage, partial coverage, and out-of-coverage scenarios to support diverse use cases, including sidelink-based positioning for V2X applications [180].
In Release 18, 3GPP further advanced NR positioning with the "Study on expanded and improved NR positioning" [181]. This release introduced multiple enhancements aimed at improving positioning performance and supporting new use cases. To enhance positioning accuracy, it proposed PRS and sounding reference signal (SRS) bandwidth aggregation for intra-band carriers and NR carrier phase measurements. In addition, Release 18 supported low-power high-accuracy positioning and positioning for RedCap UEs. Considering the needs of V2X positioning, sidelink-based positioning was also studied in [181].
For AI/ML aspect, starting with Release 18, 3GPP initiated a study on AI/ML positioning, as outlined in [182]. This study explores the potential of AI/ML to improve positioning accuracy, details the general AI/ML framework, use cases for AI/ML positioning, and evaluation metrics and common KPIs for AI/ML positioning. These developments aim to complement traditional positioning methods by leveraging data-driven models to process large datasets and provide more accurate results.
TABLE I: 5G Positioning performance requirements
Positioning Service Level Positioning requirements
Absolute Positioning Relative Positioning     Service Availability Latency Mobility
Horizontal Vertical Horizontal Vertical
Level 1 $10$ m $3$ m N/A N/A $95$ 1 s $\leq 30$ km/h (indoor) $\leq 250$ km/h (rural and urban) $\leq 500$ km/h (trains)
Level 2 $3$ m $99$
Level 3 $1$ m $2$ m
Level 5 $0.3$ m
Level 7 N/A N/A $0.2$ m $0.2$ m $\leq 30$ km/h (indoor and outdoor)
Level 4 $1$ m $2$ m N/A N/A $99.9$ $15$ms $\leq 30$ km/h indoor
Level 6 $0.3$ m $2$ m $10$ms $\leq 30$ km/h indoor
IV-B KPIs for Wireless Positioning
Positioning performance is assessed through a set of critical metrics that ensure the accuracy and reliability of location services. The 5G system is designed to deliver positioning services that meet the performance requirements outlined in Table I. These requirements are inclusive of all UE types, including specialized UEs such as V2X and machine-type communication (MTC) devices.
In 3GPP [177], 5G positioning use cases are divided into 5G positioning service area and 5G enhanced positioning service area. The former includes indoor and outdoor (rural and urban) scenarios. Indoor includes location inside buildings such as offices, hospital, industrial buildings. Outdoor scenarios support the positioning of vehicles and trains at speeds up to 250 km/h and 500 km/h. And the 5G enhanced positioning service area further supports dense urban areas (up to 60 km/h), vehicles in tunnels, and railway positioning.
In addition to horizontal and vertical positioning accuracy, 3GPP Release 19 also introduces positioning service availability and positioning service latency to support mission-critical services. A total of 7 positioning service levels are defined, including 6 levels of absolute positioning accuracy requirements and 1 relative positioning accuracy requirement. In the absolute positioning accuracy requirements, the horizontal accuracy ranges from 10 m to 0.3 m, and the vertical accuracy ranges from 3 m to 2 m. The positioning service availability ranges from 95% to 99.9%. The scenarios of positioning service levels 1, 2, 3, and 5 require a 1s level latency, while positioning service levels 4 and 6 require a ms-level latency, which is only required in indoor scenarios of 5G enhanced positioning services, such as collaboration and collision avoidance of mobile robots and factory scenarios. For the relative positioning scenario of positioning service level 7, it is applicable to the positioning of two UEs within 10 m or the distance between a UE and a 5G positioning node within 10 m.
When evaluating AI/ML based positioning, in addition to the common KPIs mentioned above, the performance of the AI model needs to be considered. For example, various overheads include open-air overhead, auxiliary information overhead, model delivery and transmission, etc. In addition, the complexity of the model’s reasoning needs to be considered, which includes the complexity of pre-processing and post-processing, the computational complexity in tera operations per second (TOPS), floating-point operations per second (FLOPS), and multiplication-accumulation operations (MACs), and the potential difference between actual complexity and evaluation complexity due to platform dependencies and optimization solutions. In addition, regardless of the underlying algorithm, model complexity should be reported in terms of the number of real-valued model parameters and operations. Finally, for model monitoring, performance indicators that need to be considered include the accuracy and relevance of monitoring indicators, related overhead, computational and memory complexity, and latency, which reflects the timeliness of monitoring results and response operations.
IV-C Advancements in AI Positioning within 3GPP
To address the ongoing challenges in 5G positioning technologies, particularly in complex environments with multipath effects and NLOS scenarios, 3GPP Release 18 introduced an AI/ML study initiative [182]. This initiative explores the potential applications of AI/ML in positioning to improve accuracy and reliability under challenging conditions. In this section, we present the 3GPP standards for AI/ML-driven positioning from three perspectives: lifecycle management (LCM) framework, model deployment, and model inputs and outputs.
IV-C1 Lifecycle Management for AI/ML models
Figure 8: Schematic diagram of AI/ML model LCM.
Taking into account the needs of AI positioning algorithms, such as data collection, model training, updating, inference, and transmission, 3GPP Release 18 defines a comprehensive LCM framework for AI/ML positioning modules. As shown in Fig. 8, the LCM framework encompasses several key stages:
  • Data Collection: The data collection function provides the input data required for model training, management, and inference. This data includes training data, monitoring data, and inference data, all essential for AI/ML model operation.
  • Model Training: The model training function executes the training, validation, and testing of AI/ML models. It may also generate performance metrics for model evaluation. Additionally, this function handles data preparation, such as preprocessing, cleaning, formatting, and transformation, using the training data provided by the data collection function.
  • Model Management: The model management function oversees the operation and monitoring of AI/ML models, including decisions related to model selection, activation, switching, and rollback. It ensures the correctness of inference operations based on data received from data collection and inference functions.
  • Model Inference: The inference function applies the AI/ML model to input data provided by the data collection function to produce outputs. This function also performs data preparation, such as preprocessing, when required, to ensure accurate inference results.
  • Model Storage: The model storage function stores trained or updated models, which can later be used for inference. This function also acts as a key point for model transfer/delivery processes and other protocol terminations.
These modules collectively address mechanisms for data collection, model training, updating, and sharing in wireless positioning, facilitating the effective deployment of AI/ML algorithms in wireless networks.
IV-C2 AI/ML Models Deployment
Figure 9: Schematic diagram of AI/ML positioning cases and categories.
Based on the role of AI in wireless positioning, AI-driven positioning techniques are categorized into two main types: Direct AI/ML Positioning and AI/ML-Assisted Positioning.
In AI/ML-assisted positioning, the AI/ML models do not directly output the UE location through end-to-end learning. Instead, they enhance the positioning process by providing improved measurements or probabilistic estimates. For example, AI/ML models can output probabilities for LOS or NLOS conditions, refine ranging estimates (such as (TOA or OTDOA), or enhance angular measurements (e.g., AOA or AOD). These outputs are used to improve the accuracy of traditional positioning techniques by addressing uncertainties and inaccuracies in the measurement process.
In contrast, direct AI/ML positioning utilizes AI/ML models to determine the UE’s location directly. These models take raw wireless channel observations as inputs and estimate the UE’s position without relying on intermediate measurement steps. A prominent example of this approach is fingerprint-based positioning, where inputs such as channel impulse response (CIR) or power delay profile (PDP) are fed into an AI/ML model to directly estimate the UE’s location.
Based on the deployment location of AI/ML models within network entities and the distinction between direct and assisted positioning methods, 5 deployment cases have been identified. As shown in Fig. 9, the AI/ML models can be deployed on the UE, gNodeB (gNB), or the location management function (LMF), enabling flexibility in their application across different network architectures and positioning:
  1. 1.
    Case 1: In this case, the AI/ML model is deployed locally on the UE. The model can support both direct AI/ML positioning and AI/ML-assisted positioning using downlink PRS.
  2. 2.
    Case 2a: In this case, the AI/ML model is deployed on the UE. The UE utilizes the AI/ML model to perform measurements on the downlink PRS and transmits the measurement results to the LMF for positioning. This scenario only supports AI/ML-assisted positioning.
  3. 3.
    Case 2b: In this case, the AI/ML model is deployed on the LMF. The UE performs measurements on the downlink PRS and transmits the results to the LMF, where the AI/ML model determines the UE’s location. This scenario only supports direct AI/ML positioning.
  4. 4.
    Case 3a: In this case, the AI/ML model is deployed on the gNB. The gNB uses the AI/ML model to perform measurements on the uplink SRS and transmits the results to the LMF for positioning. This scenario only supports AI/ML-assisted positioning.
  5. 5.
    Case 3b: In this case, the AI/ML model is deployed on the LMF. The gNB performs measurements on the uplink SRS and transmits the results to the LMF, where the AI/ML model determines the UE’s position. This scenario only supports direct AI/ML positioning.
IV-C3 Model Inputs and Outputs
For model training, training data can be generated by various network entities, including the UE, positioning reference unit (PRU), gNB, or LMF. For LMF-side model inference (Cases 2b and 3b), the input data is generated by the UE or gNB and terminates at the LMF. For gNB-side model inference (Case 3a), the input data is directly available within the gNB, reducing latency and ensuring efficient processing.
In cellular positioning, various measurements serve as critical inputs for AI/ML models to achieve accurate localization. These inputs are primarily derived from reference signals such as DL PRS and UL SRS, which are reused from existing 3GPP specifications but can be enhanced with AI-specific configurations. Key input types include high-dimensional information like delay profiles (DP), PDP, CIR, and CIR phase data. Additionally, timing-related measurements such as TOA, RSTD, RTT, and angle-related metrics extracted from CSI can also be leveraged as model inputs. Additionally, power metrics like DL PRS-RSRP and UL SRS-RSRP also contribute to improving positioning accuracy.
The outputs of AI/ML models depend on the type of positioning. For AI/ML-assisted positioning, models output refined measurements or probabilistic estimates, such as TOA, OTDOA, LOS/NLOS probabilities, or angular measurements (AOA, AOD). For direct AI/ML positioning, models output the estimated UE location.
V SOTA in AI/ML-Assisted Positioning
As discussed above, AI/ML-assisted positioning leverages AI technologies for estimating positioning-related parameters, which are subsequently used to enhance the performance of traditional positioning algorithms. This approach is necessary because, while conventional estimation techniques are effective under ideal conditions, they often face significant challenges in real-world environments, including multipath propagation, signal distortion, and hardware imperfections. The advent of ML and DL offers new possibilities for addressing these challenges by employing data-driven methods to model the inherent complexity and nonlinearity of wireless environments. ML and DL techniques not only improve the robustness of parameter estimation but also enhance real-time performance in dynamic scenarios. Based on these advancements, we have conducted a detailed survey of AI-based Positioning Parameter Estimation algorithms, with a specific focus on AI-based LOS/NLOS detection, TOA/TDOA estimation, and angle estimation algorithms.
V-A AI-based LOS/NLOS Detection
The wireless environment significantly impacts positioning accuracy. LOS scenarios, with unobstructed transmitter-receiver paths, yield more accurate distance and angle measurements, while NLOS scenarios, involving reflections and obstructions, degrade reliability due to multipath effects and delays. Thus, identifying LOS and NLOS conditions is critical for ensuring positioning accuracy.
Prior to the advancements in deep learning, significant work had already been done in this area, employing ML algorithms such as support vector machines (SVM), Random Forest, Gradient Boosting Decision Tree [183], and AdaBoost for LOS/NLOS identification. The authors in [184] compare the performance of these traditional ML algorithms in LOS/NLOS identification tasks. To address the challenges in labeled data collection, [185] applied an unsupervised ML approach called expectation maximization for Gaussian mixture models to classify LOS and NLOS components.
Recent advancements in AI technology have significantly enhanced the capability to detect these conditions. In [186], the authors compared three ML classifiers, i.e. SVM, Random Forest, and MLP, to identify LOS and NLOS in UWB systems. In [187], in vehicle-to-vehicle (V2V) scenarios, the authors used the power angular spectrum (PAS) and employed SVM, Random Forest, and ANN for LOS/NLOS classification.
Subsequently, more advanced models and algorithms have been widely studied for LOS/NLOS identification. In [188], the authors propose a CNN-based neural network using PAS. In [189], a reversible transformation method is proposed for denoising CIR data, employing CNN to identify NLOS signals. The authors in [190] propose a Morlet wavelet transform and CNN approach for NLOS identification in indoor UWB positioning. In [191], the paper proposes a LOS/NLOS identification algorithm based on one-dimensional wavelet packet analysis and CNN. To fully utilize the characteristics of wireless channels, the authors in [192] use resource block (RB)-level eigenmatrix and eigenvector, designing a neural network based on 3D CNN for NLOS detection. To reduce computational complexity, the authors in [193] use both manually extracted features and features from a CNN based on raw CIR, designing a lightweight network with only 7.39% of the computational complexity compared to traditional CNNs. Similarly, in [194], the authors design a CNN-based lightweight LOS/NLOS recognition model with CIR as input and achieved good accuracy. In [195], the authors use CNN and GRU for NLOS identification based on CIR. In substation scenarios [196], the paper designs a neural network based on self-attention to separately learn the information contained in manually extracted channel features and the original channel features to enhance identification performance. The authors in [197] propose a dual-channel neural network with the consideration of the time and time-frequency domain characteristics. In [198], the paper proposes an algorithm based on Stockwell transform and CNN, utilizing transfer learning to enhance the algorithm’s performance. Additionally, the authors in [199] propose a learning method based on CNN, Bidirectional LSTM, and transfer learning for NLOS detection. In UWB scenarios [200], the authors propose a robust NLOS identification method based on transfer learning and GAN, utilizing cross-domain mapping from source and target domains to construct representative homogeneous features for both domains. In mmWave scenarios [201], the paper models the problem as a semi-supervised anomaly detection problem by exploiting a deep autoencoding kernel density model to identify several key parameters describing sparse spatiotemporal channel responses. Additionally, the authors in [202] perform LOS/NLOS identification based on RSS and RTT instead of CIR.
Furthermore, various works have leveraged LOS/NLOS identification to improve positioning algorithms. For instance, the authors in [203] utilize LSTM to estimate approximate NLOS errors, correcting the positioning estimation results to enhance positioning accuracy. In UWB scenarios, the paper [204] converts CIR to Gramian angular fields (GAF) and uses CNN for NLOS identification, greatly enhancing ranging accuracy and reducing positioning errors based on identification results. In 5G channels, the authors in [205] use the angle-delay channel response matrixs for NLOS identification and select different positioning methods based on the identified LOS/NLOS results to enhance positioning accuracy.
V-B AI-based TOA/TDOA Estimation
As mentioned above, TOA and Time TDOA are fundamental techniques in wireless positioning, providing the basis for accurate positioning by measuring the time it takes for a signal to propagate between transmitters and receivers. However, traditional TOA/TDOA estimation methods face significant challenges in real-world environments, including the presence of multipath propagation, NLOS conditions, and hardware imperfections. Recent advancements in AI can effectively enhance TOA/TDOA estimation by leveraging data-driven learning techniques.
Early studies applied traditional ML algorithms to TOA estimation, focusing on simple models to mitigate errors in distance measurements [206, 207]. For example, the authors in [206] propose a SVM approach for error mitigation in UWB ranging systems. In [207] the authors use kernel principal component analysis for TOA estimation. In [208], the authors utilise SVM and semi-supervised learning to achieve ranging estimation. While these methods provided initial improvements, their inability to effectively interpret high-dimensional channel features limited their robustness and adaptability in complex environments.
To overcome the limitations of traditional ML, deep learning techniques have been employed to enhance TOA/TDOA estimation. These methods leverage the ability of neural networks to model complex, nonlinear relationships in data, achieving superior accuracy [209, 210, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223]. For example, in [209], the authors introduce an ANN-based RTT estimator for WiFi. The authors in [210] utilize ANN and radial basis function (RBF) networks to output distance errors rather than direct positions, enhancing TOA measurement accuracy. In addition, the authors in [211] propose error-mitigating ANNs to further refine TOA/TDOA estimates. In [212], the paper uses deep CNNs to exploit WiFi preamble structures for TOA estimation, effectively handling multipath environments. Furthermore, the paper proposes a novel dictionary filtering method leveraging neural networks for denoising and compressive sensing for channel impulse response extraction [213]. In [214], the authors use neural networks to generate high-resolution CIR for accurate TOA estimation. In the 5G system, the authors propose a CNN-based algorithm for TOA estimation [215]. In the IoT system, the authors in [224] propose a CNN-based algorithm by generating fine-grained features from full-band and resource-block-based reference signals, leveraging spectrogram-like cross-correlation feature maps with ML to directly project time-frequency domain variations into TOA results. In [216], the authors combine neural networks with Kalman Filters for TOA estimation using 5G downlink signals. In addition, the paper [217] addresses the challenge of false peaks in UL transmissions of LTE-LAA with a CNN-based TOA estimator tailored for the Block Interleaved Frequency Division Multiplexing (B-IFDM) structure. Considering hardware imperfections, the authors in [218] propose a CNN-based TOA positioning algorithm calibrated with CIR. Furthermore, considering the time dependence of CIR, the paper [221] proposes a BiRNN for accurate TOA estimation from CIR and an MLP for trilateration positioning. In addition, there are some works that consider the joint estimation of AOA and TOA to achieve higher accuracy estimation [225, 226, 227].
In addition to the fundamental approach of end-to-end training with neural networks, recent research also leverages advanced AI techniques to address challenges such as data scarcity [113, 208, 228] and robustness [229] in TOA/TDOA estimation. To solve the difficulty of obtaining datasets, in [113], the paper combines neural networks with the Fine Timing Measurement (FTM) protocol to enhance WiFi RTT ranging. This work introduces an unsupervised learning framework that uses naturally accumulated sensor data to reduce data collection overhead. The authors in [228] utilize sensor data gathered during regular application use to reconstruct device trajectories. The reconstructed data is used to design an unsupervised learning technique for TOA estimation. The authors in [208] propose a semi-supervised learning approach for NLOS identification and mitigation, leveraging self-training with unlabeled measurements to achieve a 94% identification probability and reduce ranging error by 10%. To improve the robustness of the algorithm, the authors [229] propose inter-instance variational autoencoders (IIns-VAEs), which use variational inference with latent variables to simultaneously estimate distance and identify environmental conditions, demonstrating strong performance on real-world datasets.
V-C AI-based Angle Estimation
In the era of traditional ML, algorithms such as Random Forest [230], SVM [231] and Regression Trees [232] have been widely employed for angle estimation. For example, the authors in [230] present an Random Forest-based optimization algorithm for AOA positioning, using weighted error mitigation to improve accuracy in NLOS environments. In [231], the paper presents an SVM-based AOA estimation method for vehicular communications. Although these methods are effective to some extent, they usually perform poorly in complex environments where multipath and NLOS effects dominate.
Compared to traditional ML techniques, neural networks have been employed due to their superior fitting capabilities, enabling more accurate angle estimation. Some studies demonstrate the potential of neural networks for angle estimation based on DNN [233, 234, 235, 236]. For instance, in [233], a deep learning-based framework integrates DNNs for efficient channel and DOA estimation in massive MIMO systems. To achieve super-resolution AOA estimation, the authors in [232] firstly obtain the MUSIC spectrum and then regard it as the input of the DNN, which increases the AOA estimation performance. In [236], the paper proposes a neural network-based method to reconstruct full-array covariance matrices from subarray samples, enabling improved DOA estimation with MUSIC. For device-free localization, the authors in [237] propose a learning-based AOA estimation, using a classifier to identify multipath components and a multilayer perceptron to correct AOA errors.
With the development of neural network models, the accuracy of angle estimation has been further improved based on CNN [238, 239, 240, 80, 241, 242, 243, 244, 245, 246, 247, 248, 249, 250], LSTM [251], Transformer [252, 253] and other advanced models[254]. The authors in [255] propose a deep convolutional network leveraging sparsity priors for efficient and precise DOA estimation, offering near real-time performance and improved robustness at low SNR. In [238], the paper introduces a deep ensemble learning approach for 2D DOA estimation, combining multiple independently trained CNNs to map spatial covariance matrices to azimuth and elevation angles to achieve angle estimation. In [239], the authors demonstrate the importance of phase features to improve DOA estimation accuracy, and propose a neural network framework, including DNN, 1-D CNN, and 2-D CNN with zero-padding for DOA estimation. Considering low-SNR conditions, the authors in [240] present a CNN-based method for robust DOA estimation, framing the problem as a multi-label classification task using sample covariance matrices. In [80], the authors introduce RFDOA-Net, a CNN-based DOA estimation method for UAV localization using nonuniform linear antenna arrays. In [241], the authors propose SDOA-Net, offering robustness to array imperfections and scalability to any target number with faster convergence. The authors in [242] introduce MoD-DNN, a model-driven DNN for AOA estimation, reformulating it as spatial spectrum image reconstruction, combining a CNN with a sparse conjugate gradient algorithm for automatic phase error calibration. Considering that the channels in wireless systems are complex, the authors in [243] introduce a complex-valued deep learning framework, and use virtual covariance matrices to handle spherical wave effects and complex signal features for near-field DOA estimation. Then, in [244], the paper employs a complex neural network-based deep learning approach and a parameterized algorithm for joint AOA and AOD estimation, enhancing computational efficiency via channel matrix preprocessing and coarse timing estimation. In addition, the authors in [251] propose an LSTM-based DOA estimation method that enhances phase features to improve accuracy and robustness to array imperfections. In [252], the authors propose a Transformer-based signal denoising network with temporal attention to enhance AOA estimation accuracy in indoor NLOS environments. The authors in [253] propose a Transformer-based sliding symbol detection method for ISAC, enabling simultaneous symbol detection, AOA and time delay estimation. Additionally, in [254], the paper proposes a cascaded neural network for DOA estimation of closely spaced sources, incorporating an SNR classification network and specialized estimation subnetworks for improved accuracy under low SNR and limited snapshots.
Furthermore, advancements in AI technologies have further enhanced the performance of angle estimation algorithms. Autoencoders, known for their powerful feature extraction capabilities, have been increasingly applied in this domain. For example, in [256], the paper introduces a DNN framework combining autoencoders and parallel classifiers to achieve robust DOA estimation. This framework effectively addresses array imperfections and demonstrates strong generalization to unseen scenarios. In addition, transfer learning can improve the robustness of the DOA algorithm through domain transfer. The authors in [40] propose a transfer learning method using a ResNet for AOA estimation in massive MIMO systems, leveraging shared features across channel models to reduce data requirements and avoid training separate networks for each channel. Similarly, in [227], the authors propose a deep transfer learning-based DOA and TOA joint estimation algorithm using a multi-task network with shared-private structure, enabling efficient fine-tuning across different SNR scenarios for improved accuracy and reduced complexity. In the vehicle positioning system [257], the authors perform 2D DOA estimation for incoherently distributed sources with massive MIMO, employing dual 1D-CNNs, transfer learning, and attention mechanisms for improved accuracy, robustness, and efficiency.
VI SOTA in Direct AI/ML Positioning
In this section, we discuss the SOTA in direct AI/ML positioning, categorizing techniques into Fingerprint-based Positioning, Knowledge-Assisted AI Positioning, and Channel Charting-Based Positioning based on modeling approaches and reliance on auxiliary information.
  • Fingerprint-based Positioning: This method directly applies AI/ML to map wireless signals (e.g., RSS or CSI) to user positions using pre-collected signal fingerprints. It achieves high accuracy positioning with comprehensive fingerprint databases but requires extensive data collection and frequent updates, making it resource-intensive and less adaptable to dynamic environments.
  • Knowledge-assisted AI positioning: This approach integrates wireless domain and geometric domain knowledge into AI models to enhance fingerprint-based positioning. By embedding prior knowledge, it improves learning efficiency, positioning accuracy, and generalization, reducing reliance on exhaustive datasets and performing well in complex, dynamic environments.
  • Channel charting based positioning: Channel charting based positioning uses self-supervised learning to model relationships between signal sampling points, creating a pseudo-coordinate system from raw CSI data without relying on absolute positions or detailed environment knowledge. It is scalable and reduces dependency on labeled datasets, making it suitable for dynamic and large-scale scenarios.
VI-A Fingerprint-based positioning
As discussed earlier, fingerprint-based positioning involves three critical processes: fingerprint collection and feature extraction, fingerprint database construction, and fingerprint database updating. Below, we provide a detailed introduction to these three components.
VI-A1 Fingerprint Collection and Feature Extraction
In the fingerprint collection and feature extraction stage, wireless signals are collected offline at RPs, and relevant features are extracted to serve as fingerprints. The simplest forms of fingerprints include RSS [258, 259, 260, 261, 262, 263], RSRP [264], RSRQ [265, 264], and TOA measurements [145, 146]. However, these types of fingerprints are often insufficiently detailed and lack the reliability required for positioning performance due to factors such as device errors and measurement noise.
CSI, with its higher-dimensional signal representation, captures the essential characteristics of wireless signal transmission, encapsulating comprehensive information about the wireless environment. This makes CSI a valuable resource for achieving high-precision positioning. However, the high-dimensional nature of CSI also poses challenges. First, extracting location-related features from raw CSI data is not easy because it is usually difficult to define which features are more important for the positioning task. Second, higher-dimensional CSI also increases the complexity of AI algorithms.
Therefore, signal preprocessing of CSI is crucial to effectively extract meaningful positioning features from raw CSI data. Some works directly utilize CFR [266, 267, 268] as model input. To extract features like AOA and TOA more effectively, several studies apply the discrete fourier transform (DFT) [269, 270, 271, 272] for CIR feature extraction. Furthermore, techniques such as quantization, normalization, grayscale processing, and principal component analysis [269] are employed to reduce parameter count and save storage space. For example, in [273], the paper combines RSS and CSI, using CNN-based feature fusion and gradient blending to improve accuracy. In [274], the paper proposes a positioning framework based on deep learning under commercial LTE systems, utilizing real-time CFR data collected via software-defined UE and employing a time-domain fusion approach to enhance positioning accuracy and robustness. The authors in [275] extract the energy coupling matrix in the refined beam domain as CSI-based fingerprints. In [271], DFT-extracted angle-domain channel power matrices are used as fingerprints, with spectral clustering for efficient matching and weighted K-nearest neighbors (WKNN) for location estimation. The authors in [276] propose a CSI quality control module that combines the Hampel identifier, wavelet filter, and cross-correlation detector to obtain clean and stable 5G NR CSI fingerprints, as well as a linear transformation module to mitigate the effects of sampling frequency offset and carrier frequency offset. In [277], the authors utilize an angle-delay channel power matrix (ADCPM) as a high-resolution fingerprint, complemented by a novel similarity criterion, compression, and two-stage clustering for efficient database preprocessing and improved localization performance. Similarly, the paper in [95] also uses the sparsity-enhanced ADCPM fingerprints to extract multipath characteristics and improve positioning accuracy while reducing computational complexity and noise sensitivity.
[…]
VI-A2 Fingerprint Database Construction
Traditional fingerprint positioning relies on constructing a fingerprint database and determining user locations by comparing real-time measurements with the database. Interpolation methods are widely used across time, frequency, and spatial domains to reconstruct complete radio maps for positioning [279, 280]. Model-based interpolation techniques, such as linear interpolation [281], Gaussian process regression [282], Kriging interpolation [283], and Voronoi tessellation-based interpolation method [284] have been extensively utilized for this purpose. For example, in [281], the paper explores approximating RSSI values using linear interpolation and Gaussian process regression, balancing accuracy, computational complexity, and data collection time. Advanced approaches also include multicomponent optimization and sparse recovery [285], total variation norm minimization for edge-preserving ray tracing [286], and energy-modified leverage sampling to improve matrix completion [287]. To further enhance the performance of the radio map, other methods leverage positional uncertainty [282], propagation priority [288], and graph-based signal processing incorporating NLOS conditions [289]. All of these techniques attempt to construct reliable and adaptable radio maps from the perspectives of accuracy, complexity, and data efficiency.
In addition, AI-based interpolation methods have also been widely studied. In [290], the paper proposes a novel deep learning framework for radio map construction, transforming spatial interpolation into a shadowing adjustment problem and introducing a neural network structure suitable for this task. In [291], the authors propose an LSTM-based deep learning method for radio environment map reconstruction in V2X scenarios. The authors in [292] propose a super-resolution radio environment map construction method combining Kriging interpolation, dictionary learning, and Random Forest, enabling efficient and accurate high-resolution radio environment maps with reduced computation time. Similarly, the study in[293] investigates feedforward neural networks for path loss modeling to enhance Kriging-based radio environment mapping. The authors [294] propose a two-phase learning framework integrating radio propagation models with a conditional GAN to extract global propagation patterns and local shadowing effects. Additionally, DeepREM uses U-Net and conditional GAN models to estimate the radio environment map from sparse measurements without requiring any additional information [295]. Moreover, in [296], the paper leverages a fully convolutional deep completion autoencoder to learn spatial propagation structures from prior data, significantly reducing measurement requirements while enhancing accuracy. LocUNet [297] uses an end-to-end CNN for user localization using RSS from a small number of BSs, leveraging pathloss radio map estimations for robust real-time performance without the need for specific area fingerprints.
However, the fingerprint positioning approach based on the radio environment map depends heavily on the density of RPs and the efficiency of the query algorithm. It faces challenges such as high deployment costs and prolonged query times. With the advancement of AI technologies, existing studies increasingly replace fingerprint databases with AI models. AI algorithms learn the mapping between wireless fingerprints and locations from datasets, significantly improving positioning accuracy. This enhancement in accuracy largely depends on the design of the AI models, leading to the development of effective models using architectures like FCNN [265], CNN [261, 266, 298, 95], LSTM [262], RNN [299, 300], and Transformers [301, 269]. For example, in [302], the paper proposes a hybrid indoor positioning architecture combining CNN, LSTM, and GAN models to enhance training data and improve accuracy. DeepWiPos uses an LSTM-based framework with attention modules to fuse RSS and fingerprint spatial gradients, addressing RSS instability and spatial ambiguity [303]. The paper in [298] investigates Residual Block-based CNN algorithms, using uplink beamformed CSI fingerprints with multiple spatial dimensions at both the BS and UE. Furthermore, in [301], the authors propose CrowdBERT, a transformer-based semi-supervised algorithm that leverages crowdsourced fingerprint data and employs spatial attention encoding, RSS-token masking, and fine-tuning to enhance spatial feature extraction.
To further enhance the robustness of the algorithm, some studies leverage advanced AI techniques such as contrastive learning [304], semi-supervised learning [305, 306], autoencoders[276], and meta-learning-based [307, 308] to improve positioning accuracy. In [304], the authors employ contrastive learning to capture wireless channel representations. By applying stochastic channel augmentations, they generate different views of the channel and learn representations for both micro-fading and macro-fading effects. In [305], a semi-supervised learning framework using a signal-guided masked autoencoder is proposed for high-precision positioning with limited samples. Additionally, an LSTM network is incorporated to leverage CIR’s temporal characteristics for user coordinate estimation. In [306], the authors propose a semi-supervised contrastive learning technique for massive MIMO fingerprint positioning, leveraging partially labeled pilot signal data and data augmentation to pre-train an encoder with contrastive loss, followed by fine-tuning for accurate positioning. The authors in [276] propose the iPos5G system, which first uses an unsupervised deep autoencoder network to reconstruct CSI features, then employs supervised learning to improve the RBF for optimizing the similarity calculation probability model, and finally performs positioning using an amplitude-phase probability fusion function. In [307], the paper proposes a Bayesian meta-learning-based fast adaptation approach to address the challenge of outdated samples, enabling pre-trained models to quickly adapt to new tasks with improved robustness. The authors in [264] propose 5G1M, a simplified 5G-based indoor positioning algorithm leveraging a Siamese network with Ghost modules, transfer learning, and trajectory fitting to reduce database dependence and enhance adaptability. Furthermore, in [309], the paper compares early and late fusion techniques, proposing a multi-task learning scheme to improve accuracy and efficiency while leveraging uncertainty estimation for enhanced fusion reliability.
To further reduce the cost of model construction, transfer learning and similar methods are utilized to migrate existing models and knowledge from one environment to another, thereby lowering deployment overhead. In [267], the paper proposes a meta-learning-based deep learning model for radio-based positioning, enabling improved transfer learning by separating environment-independent feature learning from environment-specific adaptation. TransLoc [41] is a heterogeneous knowledge transfer framework for fingerprint-based indoor localization. It addresses real-time environmental dynamics by refining source domain knowledge, creating a homogeneous feature space through cross-domain mapping, and utilizing a joint optimization algorithm for efficient knowledge transfer with minimal target domain samples. In [310], the study addresses cross-environment indoor positioning using a semi-supervised approach, introducing a deep neural forest combined with adversarial training to learn environment-invariant features, enabling robust localization without requiring annotations in new environments. Moreover, the authors in [272] present the cross-region fusion and fast adaptation framework for fingerprint localization in cell-free massive MIMO systems, leveraging cross-region fusion, meta-learning, and fine-tuning to enhance localization accuracy, achieve rapid deployment.
VI-A3 Fingerprint Database Updating
Fingerprint database and AI model aging also pose critical challenges in wireless fingerprint-based positioning. Environmental changes, such as infrastructure updates, furniture rearrangements, or time variations, can cause the database and models to become outdated, reducing positioning accuracy. Addressing this issue requires efficient methods for maintaining and updating the database and AI models.
Crowdsourcing methods provide an effective solution by opportunistically collecting labeled fingerprints during routine device usage. Crowdsourced fingerprint updates leverage opportunistic real-world data, such as high-confidence labels obtained near base stations or in areas with strong GPS signals [311, 312, 313], as well as user movement trajectories obtained from pedestrian dead reckoning (PDR) [314, 315, 316, 317, 318, 142, 319]. For example, the authors in [311] use multi-kernel transfer learning for integrating historical fingerprints and GPS-based opportunistic fingerprints to achieve one-meter accuracy with minimal update overhead. Based on trajectory continuity, the authors in [315] propose integrating a modified particle filter with a self-updating fingerprint database to enhance accuracy and robustness. Similarly, in [317], the paper leverages static mobile devices as RPs and a trajectory-matching algorithm to adapt the radio map to dynamic environments. In addition, the authors in [320] propose a crowdsourcing-based radio map update method using sparse representation and low-rank matrix recovery to address incomplete and noisy fingerprints. Moreover, managing computational demands and filtering crowdsourced data also pose significant challenges. In [321], the paper transforms crowdsourcing trace data into cluster and positioning spaces, enabling unsupervised online updates with dynamic replay memory. However, these crowdsourcing-based methods all rely on additional data such as GPS and PDR, and it is still challenging to accurately determine the location of each crowdsourced fingerprint without prior information.
Advanced techniques like continual learning [48, 322] and GANs [323] have also played a pivotal role in addressing database aging. In [323], the authors introduce RecTrack-GAN, an indoor tracking framework combining RNNs for human movement modeling and Conditional GANs for data recovery and fingerprint database updates, improving tracking accuracy. In [48], the authors propose an incremental learning-based fingerprint localization scheme using CSI images and a broad learning system, enabling rapid updates without retraining and achieving superior performance in real-world indoor environments. In [324], the paper proposes a continuous active learning method and uses uncertainty sampling to periodically adapt the database to environmental changes, improving accuracy and preventing performance degradation. In addition, based on transfer learning, the authors in [322] propose a system that automatically updates radio maps using transfer learning with altered AP identification and mapping space searching, significantly improving accuracy in dynamic environments. The authors in [308] present FeMLoc, a federated meta-learning framework for indoor localization, enabling swift adaptation to new environments with minimal data and reduced calibration effort. The authors in [268] introduce a semantic localization approach enhanced by multi-task deep domain adaptation and scenario adaptive learning, integrating environmental semantics into localization frameworks to address time-varying signal propagation challenges.
VI-B Knowledge-assisted AI Positioning
Traditional positioning methods, such as trilateration, triangulation, and multilateration, utilize geometric principles and channel propagation models to estimate position based on signal properties. While these geometry-based methods are effective under some conditions, they often face significant challenges in complex or dynamic environments, particularly in scenarios affected by multipath effects and NLOS conditions.
In contrast, fingerprint-based positioning uses AI to directly learn the mapping between signal characteristics (or features) and the position, achieving end-to-end positioning. However, as noted above, fingerprint positioning systems can experience performance degradation due to changes in the wireless environment. These systems also face significant difficulties in transferring across different environments or deployment scenarios, which limits their robustness and scalability.
To address the limitations of both traditional geometry-based and pure fingerprint-based positioning approaches, knowledge-assisted AI positioning has emerged. This approach combines the power of AI with domain knowledge, such as geometric and channel propagation insights, enabling positioning systems to achieve greater accuracy and reliability. Unlike purely data-driven, end-to-end methods like fingerprinting, knowledge-driven AI-assisted positioning integrates physical models and environment-specific features as a foundation for AI models. This fusion of knowledge allows AI algorithms to perform effectively, even in challenging environments where data alone may fall short.
Through this combination of AI with domain knowledge, knowledge-assisted AI positioning can adapt more robustly to environmental variability and reduce reliance on extensive, labeled datasets, which are often required for purely data-driven approaches. This integration enhances accuracy and improves generalization across various environments, making the positioning system more resilient to external changes. In this subsection, we will introduce the research related to wireless-knowledge-assisted positioning and geometric-knowledge-assisted positioning respectively.
VI-B1 Wireless-knowledge-assisted Positioning
A considerable amount of research focuses on integrating wireless domain knowledge into AI algorithms to optimize positioning systems. Combined with this wireless domain knowledge, AI algorithms can effectively solve problems such as wireless channel feature extraction [325, 326, 327] and hardware damage [328, 329], thereby improving algorithm robustness. Several studies exploit the intrinsic properties of wireless channels to enhance the robustness of AI-based positioning [325, 326, 327, 304, 330, 331]. For instance, to address frequency diversity challenges in wireless networks, the authors in [325] propose a multi-frequency fusion learning method. This approach first extracts position-related features independently from CSI at each subcarrier using a shareable method and then fuses these features to overcome the diversity challenges inherent to FDMA systems. In [326], a joint approach considers both frequency-domain and time-domain features. In the frequency domain, relative phase differences and received powers across resource blocks are combined, while in the time domain, parameters like AOA, RTT, and signal strength are incorporated to jointly improve positioning accuracy. Recognizing the high dimensionality of channel data, the authors in [327] propose a positioning neural network and reduce computational complexity by utilizing minimal descriptive features, including maximum power measurements and their temporal locations. In [330], the authors present a view-selective deep learning system, leveraging multiview training with supervised variational networks and dominant view classification to enhance localization accuracy in complex radio environments. Additionally, in distributed M-MIMO frameworks, the study [331] introduces a deep belief network to analyze RSS values derived from diffraction models for positioning.
Considering hardware impairments, the authors in [328] explore end-to-end positioning using an autoencoder architecture, effectively mitigating hardware-induced errors. In [329], the authors address challenges specific to RIS-aided mmWave systems, such as clock offsets between transmitters and receivers, impairments at the transmit and receive antenna arrays, and coupling effects among RIS elements. They propose a dictionary learning approach to calibrate these hardware impairments, significantly improving positioning performance in RIS-aided environments. These methods demonstrate the importance of accounting for hardware imperfections in achieving accurate and robust localization.
In addition, based on communication system properties, some studies explore the synergies between wireless positioning and other communication tasks, such as channel estimation [329] and CSI feedback [332]. For instance, the authors in [329] jointly optimize channel estimation and localization in RIS-aided mmWave systems. In [332], an integrated learning framework for CSI feedback and localization is proposed. This framework enables both tasks to mutually benefit each other, demonstrating that incorporating coarse positional data improves the accuracy of both CSI feedback and CSI-based localization.
VI-B2 Geometric-knowledge-assisted Positioning
Another approach is to combine AI algorithms with geometric knowledge and spatial constraints to improve positioning accuracy. This includes tasks such as NLOS detection, TOA/angle estimation, and the integration of physical map constraints to enhance the understanding of spatial relationships between transmitters and receivers. When combined with AI algorithms, geometric knowledge allows these systems to adjust for variations in wireless environments or spatial shifts. For instance, in OTDOA-based positioning [333], the authors enhance accuracy by incorporating NLOS indicator with RSTD measurements as inputs to a neural network, which effectively mitigates NLOS problem and improves the positioning accuracy. Further extending this, the authors in [334] use hybrid delay and angular measurements in a neural network-based weighted least squares (WLS) framework to enhance performance. Another study [335] employs a deep variational learning method to estimate position-related parameters such as distance, TDOA, and AOA, which are subsequently used to calculate the position. In [336], the authors propose MLLoc, a ML-based positioning system that fuses GNSS and mobile network signals, using ML for stable TOA estimation to calculate positions accurately. In [337], the authors propose a classification-to-regression ANN model, leveraging class probabilities from classification to compute final positioning coordinates.
In addition, positioning can be enhanced by utilizing environmental maps [338, 339], the correlation of user motion trajectories [340, 341], and the relationships between multiple BSs [342]. Incorporating prior physical map information, the authors in [338] propose a zero-shot learning framework for indoor localization using floor-plan images. A graph neural network is employed to model the relationships between APs and mobile devices for coarse localization, while floor-plan constraints refine positioning accuracy. The authors in [339] focus on multipath detection to reconstruct indoor environments. The study uses an ML model to predict dominant multipath components, identify virtual anchors, and build a generative channel model, significantly improving positioning accuracy under NLOS conditions. To address issues with outlier positioning estimates, the authors in [340] propose a context-aware localization technique that leverages historical trajectory information to refine the accuracy of anomalous points. In [341], the authors propose a DRL-based unsupervised wireless localization method, modeling localization as a Markov decision process and designing a reward-setting mechanism based on high RSS near APs for robust localization without retraining. In [342], the paper proposes a graph convolutional network based algorithm to model spatial relationships among multiple APs, extracting features from RSSI-based fingerprints for classification using an MLP.
VI-C Channel Charting based Positioning
As mentioned above, wireless-knowledge-assisted positioning builds upon fingerprint-based positioning by leveraging wireless knowledge to further enhance localization performance and reduce the reliance on extensive data collection. Building on this foundation, Channel charting based positioning eliminates the need for external RPsor exhaustive datasets by directly learning latent spatial relationships from CSI data, enabling relative or pseudo-positioning. This approach allows channel charting to be highly adaptable to dynamic environments while significantly reducing operational overhead, making it a promising alternative to traditional localization methods in rapidly changing wireless scenarios.
Numerous studies have investigated the potential of channel charting to enhance wireless positioning performance [343, 344, 345, 346, 166, 347, 348, 349, 160, 350]. The authors in [343] first introduce channel charting for positioning by proposing a unified Siamese network architecture for CSI-based localization. Their framework supports supervised, semi-supervised, and unsupervised scenarios, leveraging Sammon’s mapping extension and side information to achieve accurate positioning in both LOS and NLOS channels with minimal CSI measurements. Building on this, graph-based approaches are further applied to channel charting and wireless positioning. For example, the authors in [344] propose a semi-supervised graph-based channel charting framework for 5G localization that utilizes distributed CSI, side-information, and constrained manifold learning to construct a 2D channel chart, achieving 5.6 m localization accuracy with minimal labeled samples. In [345], a comparative study between classical model-based localization methods and channel charting highlight the limitations of channel charting in achieving global geometric accuracy. To address this, augmented channel charting is proposed by integrating model-based localization information into channel charting training, thereby improving overall performance and surpassing classical methods on evaluated datasets. Leveraging temporal correlations during sampling, the authors in [346] introduce a reference-free channel charting framework that incorporates velocity information and topological maps to transform relative charts into real-world coordinates. Further enhancements are made in [166], where a novel dissimilarity metric is introduced, incorporating angular-domain information and a deep learning-based metric. Additionally, metric fusion is proposed to integrate temporal and CSI similarities, demonstrating superior performance in sub-6 GHz massive MIMO scenarios, even under NLOS conditions. In [347], a Doppler-based loss function for channel charting is introduced, requiring only frequency synchronization to enable localization with minimal assumptions. The authors in [349] develop a geodesic distance-based metric for channel charting using synchronized CSI measurements, utilizing a Siamese network to learn global geometry for localization, which outperform traditional fingerprinting methods in real-world 5G and UWB systems. Additionally, in [160], the paper presents a multi-point channel charting approach for multi-gateway LoRa networks, applying t-SNE and k-means clustering on received power vectors to map spatial geometry and improve IoT device localization. Lastly, privacy concerns in channel charting are systematically examined in [350], focusing on user and vendor privacy risks associated with pseudo-locations and raw CSI exposure.
VII Datasets for Wireless Positioning
In recent years, several datasets have been proposed in the field of wireless positioning, including the xG-Loc dataset [351], DeepMIMO dataset [352], Wireless AI Research Dataset (WAIR-D) [353], DataAI-6G dataset [354] and the ViWi dataset [355]. However, the above datasets do not conform to the clustered delay line (CDL) channel model [356] recommended by 3GPP. Since CDL channel model is more suitable for link-level and system-level simulation and for the comprehensiveness of the study, we also examine datasets based on CDL channel model.
VII-1 xG-Loc
The xG-Loc dataset [351] is the first open dataset explicitly designed for localization algorithms and services, fully compliant with 3GPP technical reports and specifications. This dataset provides a standardized and comprehensive resource for evaluating and benchmarking localization solutions across diverse scenarios. xG-Loc is structured into 28 compressed directories, representing 28 unique configurations of 3GPP-standardized scenarios, bandwidths, and central frequencies. These configurations span multiple frequency ranges, including FR1 (microwaves), FR2 (millimeter waves), and FR3 (upper mid-band), thereby enabling comprehensive analysis across various operational conditions. The scenarios considered in the xG-Loc dataset align with 3GPP 38.901 [356].
Each directory contains a rich set of data, including PRS and SRS, measurement data, analytics, position estimates, and channel quality indicators (CQIs). The dataset is organized in text and JSON files to facilitate flexible usage and compatibility with various ML frameworks and localization algorithms.
VII-2 5G-NR-data-generator
The 5G-NR-data-generator [357] is a generalized channel dataset generator that adopts the CDL channel model of the 5G NR standard [356]. This generator allows users to customize various channel parameters to meet specific needs and supports the generation of MIMO channels.
Figure 10: Dataset generation process of 5G-NR-data-generator.
The data generation process is shown in Fig. 10. The configurable channel parameters $S^{\prime}$ include a variety of settings, such as the number of RBs, subcarriers, bandwidth, the number of BS and UE antennas, and the channel center frequency. Based on the user-defined parameters $S^{\prime}$ or $S_{u}^{\prime}$, the 5G-NR-data-generator performs ray-tracing simulations to model the propagation environment. The ray-tracing simulation determines the propagation paths from the BS to the UE by utilizing the geometric surface data provided in the map files.
These map files and terrain data are sourced from the open-source Open Street Map (OSM) project [358]. Using the terrain and map data, the simulation calculates key channel characteristics, including propagation delays, AOA, AOD, LOS or NLOS labels, and path gains. These outputs are then combined to construct the final channel model.
VII-3 DeepMIMO
Figure 11: Dataset generation process of DeepMIMO.
DeepMIMO [352], developed by researchers from Arizona State University, is a versatile and generic deep learning dataset designed specifically for mmWave and massive MIMO applications. The dataset provides a comprehensive framework for generating channel datasets tailored to various tasks in wireless communication.
The channels in the DeepMIMO dataset are constructed using accurate ray-tracing data obtained from the Wireless InSite simulator [359]. This ray-tracing simulation captures the dependence of the wireless channels on the geometry and materials of the environment, as well as the spatial locations of the transmitter and receiver. This feature is crucial for ML applications in mmWave and massive MIMO systems, as it ensures the dataset realistically reflects the physical characteristics of the propagation environment.
Based on DeepMIMO, researchers can customize the dataset by configuring a set of parameters to suit their specific ML tasks. These parameters control various system and channel elements, including the number of antennas, subcarriers, channel paths, and other factors. Therefore, the DeepMIMO dataset is defined by two primary components: a scenario and a set of customizable parameters. The process of using the DeepMIMO dataset generation framework is straightforward, as illustrated in Fig. 11.
VII-4 WAIR-D
The WAIR-D [353] is developed by researchers from Huawei and Zhejiang University to provide a wireless dataset that replicates various environments closely resembling real-world conditions. The dataset is generated using a 3D ray-tracing simulator, PyLayers [360], in conjunction with the OSM [358]. PyLayers utilizes environmental information, such as maps containing the locations of BSs and UEs, to calculate radio propagation paths for each radio link between BSs and UEs. The use of OSM provides details about real-world building layouts and street directions, enhancing the authenticity and diversity of the generated data.
WAIR-D offers flexibility in generating channels with customizable wireless system parameters, supporting five carrier frequencies, i.e., 2.6 GHz, 6 GHz, 28 GHz, 60 GHz, and 100 GHz, enabling applications in both sub-6 GHz and mmWave scenarios. Users can configure system bandwidth, numerology, and the number of antennas, tailoring the dataset to specific requirements. WAIR-D includes two scenarios: one with 10,000 environments featuring sparsely distributed UEs for wide coverage and another with 100 environments focusing on densely deployed UEs for high-density urban settings.
The dataset provides a rich array of details about the propagation paths and the environment, including path delays, AOA, AOD, and other channel characteristics. The data is stored in specific file formats to facilitate its use in various wireless research applications.
VII-5 DataAI-6G
Figure 12: Dataset generation process of the ViWi dataset.
The DataAI-6G dataset [354], developed by researchers from Beijing University of Posts and Telecommunications and the China Mobile Research Institution, is specifically designed for AI-6G research. A key feature of this dataset, and its major advantage over the DeepMIMO and WAIR-D datasets, is its incorporation of spatial non-stationary features and high-mobility characteristics. With the adoption of ultra-massive MIMO and the increasing demand for high-mobility communications, these features are becoming more prominent, making it essential to include them in the dataset. This inclusion significantly enhances positioning accuracy in corresponding scenarios.
A notable advantage of the DataAI-6G dataset [354] is its generic framework, which allows researchers to configure parameters according to their specific requirements. This framework enables users to input raw channel parameters and generate customized datasets tailored to their research needs. The dataset provides detailed channel information obtained from the Wireless InSite ray-tracing simulator [359], including parameters such as AOD, AOA, delay, phase, power of each path, path loss between each pair of transceiver antennas, and features such as spatial non-stationarity and mobility.
Therefore, the dataset is particularly well-suited for validating mobility in cross-band scenarios, making it highly applicable to advanced 6G research. Furthermore, it is fully compatible with ML and AI algorithms, enabling researchers to explore innovative solutions for mobility and non-stationary environments in 6G systems.
VII-6 ViWi
The ViWi dataset [355] is designed specifically for vision-aided wireless communications research. It serves as a parametric, systematic, and scalable data generation framework, leveraging advanced 3D modeling and ray-tracing software to produce high-fidelity synthetic wireless and vision data samples for identical scenes. By combining vision and wireless data, the dataset facilitates research at the intersection of these domains, supporting innovative approaches to wireless communication and positioning.
The ViWi dataset includes diverse signal characteristics such as AOD, AOA, path gains, and images. It also provides detailed user location data, enabling researchers to explore localization and communication techniques in visually rich environments. The dataset demonstrates flexibility and diversity through its parametric customization capabilities. Visual raw data is stored in JPEG and MAT formats, while wireless raw data comprises three MAT files per transmitter, providing a comprehensive and well-structured resource for research. The dataset generation process consists of three stages, as illustrated in Fig. 12:
  • Scenario Definition: Defines physical layout, transmitter and receiver locations, and environmental parameters.
  • Raw-Data Generation: Utilizes 3D modeling and ray-tracing tools to produce realistic wireless and visual data.
  • Parametrized Processing: Processes raw data into structured formats tailored to research needs.
VIII Challenges and Opportunities
Wireless AI positioning presents both significant challenges and exciting opportunities. This section explores the key challenges and potential opportunities in this field.
VIII-A Challenges in AI-driven Wireless Positioning
Wireless AI positioning has witnessed remarkable advancements, yet several challenges hinder its widespread adoption and optimal performance. These challenges stem from the intrinsic complexities of wireless environments, computational constraints, and the integration of emerging technologies. Below, we outline the key challenges.
VIII-A1 Data Collection
The challenges in data collection for AI-driven wireless positioning are multifaceted, stemming from the dynamic, diverse, and sensitive nature of wireless data. High-quality labeled datasets are essential but difficult to obtain due to the diverse propagation characteristics across environments (urban, rural, indoor, outdoor) and the dynamic nature of wireless signals [361]. Datasets quickly become outdated, necessitating frequent updates, while over-reliance on synthetic datasets often fails to capture real-world complexities, leading to reduced model robustness. Furthermore, the discrepancies between simulated and real-world channel characteristics make it difficult for models trained on synthetic data to perform effectively in practical environments, significantly increasing the workload associated with real-world data collection. Additionally, the massive volume and high dimensionality of wireless data, particularly from advanced technologies like massive MIMO and mmWave, pose computational challenges. Outliers and noise in raw data further degrade model performance, requiring robust preprocessing and anomaly detection techniques [362].
VIII-A2 Accuracy in Complex Environments
Achieving high positioning accuracy in complex environments remains a significant challenge for AI-driven wireless positioning systems. Real-world scenarios such as urban areas, highways, and indoor settings are characterized by dynamic and unpredictable factors that reduce the precision of positioning algorithms. These factors include multipath propagation, NLOS conditions, high-speed mobility, interference, and environmental obstacles. Under these circumstances, signals are reflected, refracted, and scattered, leading to complex and distorted channel distributions. While AI models are capable of learning intricate patterns, they require extensive and diverse data to generalize effectively in such environments. However, collecting such data is often impractical or infeasible. Consequently, the challenge of achieving reliable accuracy in complex and dynamic environments continues to impede the practical deployment of AI-driven wireless positioning systems.
VIII-A3 Model Generalization
As with most AI applications, one of the key challenges in AI-driven wireless positioning is ensuring model generalization across diverse environments and deployment scenarios. Variations in signal propagation characteristics, device configurations, and environmental dynamics often cause AI models trained under specific conditions to perform poorly when applied to new or unseen environments. For instance, a model trained in an urban setting may struggle in rural or indoor environments where channel characteristics differ significantly. While advanced techniques such as domain adaptation, transfer learning, and multi-task learning can partially address this issue, their ability to fully resolve the complexity and dynamic nature of wireless positioning environments remains limited.
VIII-A4 Resource Constraints
Resource constraints pose a significant barrier to the deployment of AI-based positioning systems, particularly on edge devices such as smartphones, IoT sensors, and UAVs. AI model inference and training typically demand substantial computational power, memory, and energy, which often exceed the capabilities of these devices. Additionally, applications requiring real-time positioning, such as autonomous driving, face added complexity when processing high-dimensional data from advanced technologies like massive MIMO and mmWave systems. Developing lightweight models and efficient algorithms that balance accuracy and computational efficiency is critical but remains a challenging endeavor.
VIII-A5 Integration with Emerging Wireless Technologies
The integration of AI-driven positioning systems with emerging wireless technologies such as RIS, massive MIMO, THz communications, and NTN introduces significant challenges. These technologies offer unique capabilities, such as enhanced spatial resolution and broad coverage, which can greatly improve positioning accuracy. However, leveraging these advantages requires the development of new AI models and algorithms, adding complexity to their seamless integration with positioning systems. Additionally, designing scalable AI models capable of adapting to the diverse characteristics of different wireless technologies remains a substantial challenge, further complicating their deployment in real-world scenarios.
VIII-A6 Security and Privacy Concerns
Security and privacy are critical concerns in AI-driven wireless positioning. Raw data used for positioning, such as CSI, often contains sensitive user information, including precise locations and device identifiers, making it vulnerable to misuse, tampering, or leakage during storage and transmission. Positioning systems are also susceptible to attacks like spoofing, interference, and eavesdropping, which can compromise their reliability and accuracy. Additionally, the privacy of AI models themselves must be safeguarded. Privacy-preserving techniques such as federated learning and differential privacy offer promising solutions, but their deployment introduces additional computational overhead and complexity. Striking a balance between security and privacy requirements and maintaining system performance remains an ongoing challenge.
VIII-B Opportunities in AI-Driven Wireless Positioning
While wireless AI positioning faces numerous challenges, it also presents unprecedented opportunities for innovation and application across various domains. These opportunities arise from advancements in AI technologies, the evolution of wireless networks, and the growing demand for high-precision positioning solutions.
VIII-B1 Enhanced Accuracy Through Advanced AI Technologies
In recent years, AI technologies have advanced rapidly, presenting immense potential for overcoming the limitations of traditional wireless positioning. Cutting-edge AI models and techniques, such as Transformers, self-supervised learning, and large language models (LLMs), are continuously emerging. These innovations enable AI systems to model the nonlinear dynamics of signal propagation, achieving centimeter-level precision in challenging scenarios. For instance, Transformers can be employed to capture spatiotemporal features of wireless channels in the time and frequency domains, while graph neural networks enhance model scalability and adaptability. Moreover, self-supervised learning methods allow AI systems to extract latent representations from unlabeled datasets, creating robust models even in the absence of extensive labeled data. With the advancement of AI technology, advanced AI algorithms will have great potential to gradually solve various problems encountered in wireless positioning.
VIII-B2 Multimodal Data Fusion for Robust Systems
Advancements in LLMs have unlocked transformative opportunities for AI-driven wireless positioning by enabling the fusion of multiple data modalities. By integrating wireless signal data with additional inputs such as visual data, inertial measurements, and environmental context, AI-driven systems can achieve exceptional robustness and adaptability. For instance, combining CSI with vision-based inputs can significantly improve positioning accuracy in indoor environments. Furthermore, the fusion of LLM technology with wireless signal data introduces novel capabilities for wireless networks, such as enabling natural language-based navigation and query processing applications. This multimodal approach not only enhances system resilience in complex scenarios but also expands the application scope of AI-driven wireless positioning systems.
VIII-B3 Advances in Emerging Wireless Technologies
Emerging wireless technologies such as massive MIMO, RIS, NTN, and THz communications open up new research directions for AI-driven positioning. These technologies introduce novel features to wireless networks that create exciting opportunities for enhanced localization capabilities. For example, RIS offers programmable propagation environments, enabling dynamic control over signal paths, while massive MIMO arrays provide exceptionally high spatial resolution. These advancements complement AI algorithms, allowing them to leverage these unique characteristics to deliver superior positioning accuracy.
VIII-B4 Resource-Efficient AI Models
The development of lightweight and resource-efficient AI models offers a promising pathway for the widespread deployment of wireless positioning systems on resource-constrained devices. Techniques such as model pruning, quantization, and knowledge distillation enable the creation of compact AI models that reduce computational and energy requirements while maintaining high accuracy. Achieving precise positioning with small-scale models remains a crucial and impactful research challenge. Addressing this can facilitate the integration of AI-driven positioning systems into IoT devices, wearable technologies, and mobile platforms.
VIII-B5 ISAC-Enhanced AI Positioning
integrated sensing and communication (ISAC) systems [363] enhance positioning by utilizing echo signals to sense the surrounding environment and extract key parameters such as AoA, ToA, and Doppler shift. This capability provides a deeper understanding of the network environment, such as NLOS detection [364]. By incorporating sensing signals alongside communication-based positioning methods, ISAC systems can significantly enhance the accuracy and robustness of wireless positioning. In such a complex system, AI algorithms play a pivotal role by leveraging their advanced feature extraction capabilities to process the multimodal data from ISAC systems, enabling high-precision positioning even in challenging scenarios.
IX Conclusion
In this work, we primarily explores the potential of AI-driven wireless positioning technologies from the perspective of integrating AI with wireless positioning. While the focus is on cellular positioning scenarios, the study also incorporates insights from WiFi, Bluetooth, and UWB positioning to enhance the comprehensiveness of the algorithmic understanding. Specifically, we introduces the foundational knowledge of AI technologies and wireless positioning techniques, followed by a summary of 3GPP standards related to positioning and AI advancements. Subsequently, we also reviews the SOTA research in both AI/ML-assisted positioning and direct AI/ML positioning, as well as datasets commonly used for wireless positioning. Finally, we summarize the challenges and opportunities in AI-driven wireless positioning. With the continued advancement of AI technologies, the integration of AI and positioning is expected to deepen. This review aims to inspire researchers in both industry and academia, contributing to the advancement of this promising field.